### v1

Abstract
Fejér monotonicity is a well-established property commonly observed in sequences generated by optimization algorithms. In this paper, we introduce an extension of this property, called Fejér* monotonicity, which was initially proposed in [SIAM J. Optim., 34(3), 2535–2556 (2024)]. We discuss and build upon the concept by exploring its behavior within Hilbert spaces, presenting an illustrative example and insightful results regarding weak and strong convergence. We also compare Fejér* monotonicity with other weak notions of Fejér-like monotonicity, to better establish the role of Fejér* monotonicity in optimization algorithms.
Keywords:
Fejér monotonicity convex analysis optimization algorithms
MSC:
49M27 65K05 65B99 90C25
journal: JOTA
1 Introduction
A useful concept for analyzing the convergence of optimization algorithms is the notion of Fejér monotonicity. A sequence $(x^{k})_{k\in\mathds{N}}$ defined in a Hilbert space $\mathcal{H}$ is said to be Fejér monotone with respect to a nonempty set $M\subset\mathcal{H}$ if for all $x\in M$ and all $k\in\mathds{N}$ we have $\lVert x^{k+1}-x\rVert\leq\lVert x^{k}-x\rVert$. It is well-known that if $(x^{k})_{k\in\mathds{N}}$ is Fejér monotone with respect to $M$, then $(x^{k})_{k\in\mathds{N}}$ is bounded, and if all its weak cluster points belong to $M$, then the entire sequence $(x^{k})_{k\in\mathds{N}}$ converges weakly to some point lying in $M$; see, for instance, (Bauschke:1996, , Theorem 2.16), and Combettes:2009 for further details and historical notes.
The notion of Fejér monotonicity is named after Hungarian mathematician Lipót Fejér (1880-1959) who introduced this notion in Fejer:1922 and further developed in Fejer:1937 and later in Agmon:1954 ; Motzkin:1954 . Fejér was the force behind the highly successful Hungarian school of analysis. His influence on the development of mathematics in the last century should be acknowledged. Indeed, Fejér was the advisor of mathematicians such as John von Neumann, Paul Erdös, George Pólya, Pál Turán and Gábor Szegö.
Weaker notions of Fejér monotonicity have been considered in the literature. For instance, the concept of quasi-Fejér monotonicity, first presented in Ermolev:1969 and further elaborated in several works such as Iusem:1994 ; Combettes:2001a .
In this work, we explore a slightly different notion, called Fejér* monotonicity. It was first introduced inBehling:2024c as a key ingredient to show that the Circumcentered-Reflection Method (CRM) Behling:2024b , proposed by these authors, solves the Convex Feasibility Problem (CFP). The main difference with the usual Fejér monotonicity notion relies on the fact that now the nonincreasing distance property holds for the tail of the sequence, starting at some index which depends on the considered point $x\in M$.
This paper is organized as follows. We start in Section 2 with a brief discussion on the role of Fejér monotonicity in optimization algorithms. In Section 3 we present some basic properties of Fejér* monotonicity, and we make a discussion, upon an illustrative example on properties of Fejér* monotonicity. In Section 4 we compare Fejér* sequences and quasi-Fejér sequences. Finally, in Section 5 we present some concluding remarks.
2 Fejér* monotonicity in optimization algorithms
First, let us formally introduce the notion of Fejér* monotonicity.
Definition 1 (Fejér* monotonicity)
Let $M\subset\mathcal{H}$ be a nonempty set and consider a sequence $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$. We say that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $M$ if for any point $x\in M$, there exists $N(x)\in\mathds{N}$ such that, for all $k\geq N(x)$,
$\left\lVert x^{k+1}-x\right\rVert\leq\left\lVert x^{k}-x\right\rVert.$
We address now the role of Fejér and Fejér monotonicity in connection with fixed-point and feasibility problems. Consider an operator $T:\mathds{R}^{n}\to\mathds{R}^{n}$. We remind that $T$ is a contraction if there exists $\tau\in(0,1)$ such that
$\left\lVert T(x)-T(y)\right\rVert\leq\tau\left\lVert x-y\right\rVert$
for all $x,y\in\mathds{R}^{n}$.
It is well known, and elementary, that a contraction has a unique fixed point (i.e., a point $\bar{x}\in\mathds{R}^{n}$ such that $T(\bar{x})=\bar{x}$) and that the sequence obtained by iterating the operator (i.e., the sequence $(x^{k})_{k\in\mathds{N}}\subset\mathds{R}^{n}$ defined as $x^{k+1}=T(x^{k})$ starting at any $x^{0}\in\mathds{R}^{n}$), converges to such unique fixed point.
The situation changes when we consider notions which are weaker than the contraction property. The most basic one is that of nonexpansiveness. An operator $T$ is nonexpansive if
$\left\lVert T(x)-T(y)\right\rVert\leq\left\lVert x-y\right\rVert$
for all $x,y\in\mathds{R}^{n}$.
The set of fixed points of a nonexpansive operator $T$ may be empty, or infinite, and the behavior of the sequence $(x^{k})_{k\in\mathds{N}}$ obtained by iterating $T$ may fail to converge, but given any fixed point $\bar{x}$ of $T$, it follows immediately from Section 2 that $\left\lVert x^{k+1}-\bar{x}\right\rVert\leq\left\lVert x^{k}-\bar{x}\right\rVert$, i.e., the sequence is Fejér monotone to the set $M$ of fixed points of $T$, and thus it enjoys the basic properties of Fejér monotone sequences, namely that it is bounded, and that if some of its cluster points belongs to $M$, then the sequence $(x^{k})_{k\in\mathds{N}}$ converges to it; see, for instance, Bauschke:2017a for more on nonexpansiveness and Fejér monotonicity.
Among the nonexpansive operators, projections onto convex sets are quite significant. We remind that, given a closed and convex set $C\subset\mathds{R}^{n}$, the projection onto $C$, $P_{C}:\mathds{R}^{n}\to C$, is defined as $P_{C}(x)\coloneqq\operatorname*{arg\,min}_{y\in C}\left\lVert x-y\right\rVert$. Convex combinations of projections (i.e., operators of the form $\bar{P}=\sum_{i+1}^{m}\alpha_{i}P_{C_{i}}$ with $\alpha_{i}\in[0,1],\sum_{i=1}^{m}\alpha_{i}=1$, and $C_{i}\subset\mathds{R}^{n}$ convex) are nonexpansive, and the same holds for compositions of projections, i.e., operator of the form $\widehat{P}=P_{C_{m}}\circ\dots\circ P_{C_{1}}$. The set of fixed points of a projection $P_{C}$ is $C$, and the sets of fixed points of $\bar{P}$ and $\widehat{P}$, as defined above are equal to the set $C:=\cap_{i=1}^{m}C_{i}$. Iterating the operators $\bar{P}$ and $\widehat{P}$, one obtains the so-called Simultaneous Projections and Sequential Projections algorithms, which aim to solve the (CFP), consisting of finding a point in the intersection $C$ of closed and convex sets $C_{1},\dots,C_{m}$. See Bauschke:1996 for a detailed discussion on these methods, which can be traced back to the seminal works of Cimmino Cimmino:1938 and Kaczmark Kaczmarz:1937 . In view of the discussion above, the sequences generated by these two methods are Fejér monotone to $C$.
In order to show that the sequences generated by these algorithms do converge to a point in $C$, one needs a bit more than nonexpansiveness. Two suitable properties, slightly stronger than nonexpansiveness but weaker than the contraction property, are
  1. (i)
    firm nonexpansiveness, meaning that
    $\left\lVert T(x)-T(y)\right\rVert^{2}\leq\left\lVert x-y\right\rVert^{2}-\left\lVert(T(x)-T(y))-(x-y)\right\rVert^{2}$
    for all $x,y\in\mathds{R}^{n}$; and
  2. (ii)
    nonexpansiveness plus, meaning that
    $\left\lVert T(x)-T(y)\right\rVert\leq\left\lVert x-y\right\rVert$
    for all $x,y\in\mathds{R}^{n}$, and $T(x)-T(y)=x-y$ whenever equality holds in (ii).
Projections onto convex sets, as well as the above defined operators $\bar{P}$ and $\widehat{P}$, are firmly nonexpansive. Convex combinations of firmly nonexpansive operators may fail to be firmly nonexpansive, but they are nonexpansive plus (see Arefidamghani:2023 ). With the help of the Fejér convergence properties and the notion of firm nonexpansiveness, it is easy to prove that the Simultaneous and the Successive Projection algorithms do converge to a point in $C$, whenever $C$ is nonempty (see Bauschke:1996 ).
One way to accelerate these methods for CFP is to approximate the set $C$ from the inside, i.e., to replace at iteration $k$ the set $C$ by a set $C^{k}$ such that
$C^{k}\subset C^{k+1}\subset\dots\subset C,$
for all $k$. For instance, if $C\coloneqq\{x\in\mathds{R}^{n}\mid g(x)\leq 0\}$ for some convex $g:\mathds{R}^{n}\to\mathds{R}$, then $C^{k}$ could be chosen as
$C^{k}\coloneqq\{x\in\mathds{R}^{n}\mid g(x)+\epsilon_{k}\leq 0\},$
where $(\epsilon_{k})_{k\in\mathds{N}}$ is a sequence of positive real numbers decreasing to 00. Then one could consider the sequence $(x^{k})_{k\in\mathds{N}}$ given by $x^{k+1}=T_{k}(x^{k})$, where $T_{k}$ is the projection onto a convex set which contains $C^{k}$. In such a case, the nonexpansiveness of $T^{k}$ gives $\left\lVert x^{k+1}-y\right\rVert\leq\left\lVert x^{k}-y\right\rVert$ but only for points $y\in C^{k}$, not in the whole set $C$.
However, if we take $C^{k}$ as in Section 2, and $C$ has nonempty interior, any point $y\in$ int$(C)$ (i.e., such that $g(y)<0$), will belong to $C^{k}$ for all $k$ bigger than $k(y)$, defined as the first integer such that $\epsilon_{k(y)}<-g(y)$, due to Section 2, but this is exactly the same as saying that $(x^{k})_{k\in\mathds{N}}$ is Fejér monotone to int$(C)$.
Hence, Fejér* monotonicity is precisely what is needed for the analysis of algorithms which use approximation schemes as in Section 2. Among them, we mention the algorithms in Iusem:1986a ; Iusem:1986 for CFP. These methods employ the scheme in Section 2 combined with operators related to $\bar{P}$ and $\widehat{P}$, where the projections onto the sets $C_{i}$, approximated from the inside by sets $C^{k}_{i}$, as in Section 2, are replaced by projections onto hyperplanes separating $x^{k}$ from $C^{k}_{i}$. In these cases, the Fejér monotonicity ensures convergence to a point lying possibly in the boundary of $C$ (despite the fact that the sequence is Fejér monotone to int$(C)$). A similar situation occurs with the algorithm in Behling:2024c , where the algorithm in Iusem:1986a is enhanced through the addition of a circumcentering step, using the machinery developed during the last few years in Araujo:2022 ; Arefidamghani:2021 ; Arefidamghani:2023 ; Behling:2018 ; Behling:2018a ; Behling:2020 ; Behling:2021b ; Behling:2023 ; Behling:2024 ; Behling:2024b ; Behling:2024c .
It is worthwhile to note that in the case of Iusem:1986 ; Iusem:1986a ; Behling:2024c , under a Slater hypothesis (i.e., when $\operatorname{int}(C)$ is nonempty), if the $\epsilon_{k}$’s decrease to 00 slowly enough, e.g., so that $\sum_{k=0}^{\infty}\epsilon_{k}=+\infty$, for instance $\epsilon_{k}=1/k$, convergence is indeed finite, i.e., there exists $\bar{k}$ such that $x^{\bar{k}}$ belongs to $C=\bigcap_{i=1}^{m}C_{i}$.
3 Properties of Fejér* monotonicity
The first result of this section was already proved in (Behling:2024c, , Prop. 2.5). However, for the sake of completeness, we present it here.
Proposition 1
Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be a Fejér* monotone w.r.t. a nonempty set $M$ in $\mathcal{H}$. Then,
  1. (i)
    $(x^{k})_{k\in\mathds{N}}$ is bounded;
  2. (ii)
    for every $x\in M$, the scalar sequence $(\left\lVert x^{k}-x\right\rVert)_{k\in\mathds{N}}$ converges;
  3. (iii)
    $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $\operatorname{conv}(M)$.
Proof
For proving (i), take any point $x\in M$. From the definition of Fejér* monotonicity, we conclude that $x^{k}$ belongs to the ball with center at $x$ and radius $\left\lVert x^{N(x)}-x\right\rVert$ for all $k\geq N(x)$. Consequently, $(x^{k})_{k\in\mathds{N}}$ is bounded.
Item (ii) is a direct consequence of Definition 1, since the $N(x)$-tail of sequence $(\left\lVert x^{k}-x\right\rVert)_{k\in\mathds{N}}$ is monotone and bounded, so the sequence converges.
For item (iii), take any $w\in\operatorname{conv}(M)$. Thus, $w$ can be written as $w=\sum_{i=1}^{p}\lambda_{i}w^{i}$, where $w^{i}\in M$, $\lambda_{i}\in[0,1]$ for all $i=1,2,\ldots,p$, $\sum_{i=1}^{p}\lambda_{i}=1$ and $p\in\mathds{N}$. Taking into account that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $M$, for each $i=1,2,\ldots,p$, there exists $N(w^{i})$ such that $\left\lVert x^{k+1}-w^{i}\right\rVert\leq\left\lVert x^{k}-w^{i}\right\rVert$ for all $k\geq N(w^{i})$. Squaring both sides of the last inequality and rearranging the terms, we get
00 0 $\leq\left\lVert{x^{k}}\right\rVert^{2}-\left\lVert{x^{k+1}}\right\rVert^{2}-2\left\langle x^{k}-x^{k+1},w^{i}\right\rangle.$ (1)
The above inequality also holds for all $k\geq K\coloneqq\max\{N(w^{1}),N(w^{2}),\ldots,N(w^{p})\}$. Multiplying Eq. 1, for each $i=1,2,\ldots,p$, by the respective $\lambda_{i}$ and adding up, we obtain
00 0 $\leq\left\lVert{x^{k}}\right\rVert^{2}-\left\lVert{x^{k+1}}\right\rVert^{2}-2\left\langle x^{k}-x^{k+1},\sum_{i=1}^{p}\lambda_{i}w^{i}\right\rangle$ (2)
$=\left\lVert{x^{k}}\right\rVert^{2}-\left\lVert{x^{k+1}}\right\rVert^{2}-2\left\langle x^{k}-x^{k+1},w\right\rangle,$ (3)
and therefore, $\left\lVert x^{k+1}-w\right\rVert\leq\left\lVert x^{k}-w\right\rVert$, for all $k\geq K$. So $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $\operatorname{conv}(M)$, as the proof is complete. ∎
We remark that a Fejér monotone sequence with respect to a nonempty set $M\subset\mathcal{H}$ is also Fejér monotone with respect to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, the closure of $\operatorname{conv}(M)$; see (Bauschke:2023, , Lem. 2.1(ii)). Furthermore, Proposition 1(iii) guarantees that a Fejér* monotone sequence w.r.t. $M$ is as well Fejér* monotone w.r.t. $\operatorname{conv}(M)$. However, the Fejér* monotonicity cannot always be extended to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, as shown in the next example.
Example 1
Let $M\coloneqq(w^{k})_{k\in\mathds{N}}\subset\mathds{R}^{2}$ with $w^{k}\coloneqq(1/2^{k},0)$, for all $k\in\mathds{N}$. Define the sequence $(x^{k})_{k\in\mathds{N}}\subset\mathds{R}^{2}$ by $x^{0}\coloneqq(0,2)$ and, for $\ell\in\mathds{N}$,
$x^{2\ell+1}\coloneqq x^{2\ell}+(1/2^{\ell},0)\;,\quad x^{2\ell+2}\coloneqq\left(0,\sqrt{\|x^{2\ell+1}-(1,0)\|^{2}-1}\right).$
If $k$ is even, $x^{k+1}$ is obtained by moving $1/2^{(k/2)}$ horizontally to the right. If $k$ is odd, $x^{k+1}$ is the intersection between the vertical axis and the arc centered in $w^{0}=(1,0)$ passing through $x^{k}$; see Figure 1.
Figure 1: From Example 1: set $M$, first nine points of sequence $(x^{k})_{k\in\mathds{N}}$, $\operatorname{conv}(M)=(0,1]\times\{0\}$, $\operatorname{cl}\left({\operatorname{conv}(M)}\right)\setminus\operatorname{conv}(M)=\{(0,0)\}$ and limit point $x^{*}$ of sequence $(x^{k})_{k\in\mathds{N}}$.
We now show that $(x^{k})_{k\in\mathds{N}}$ is Fejér* with respect to $M$. Denote $x^{k}=(\alpha_{k},\beta_{k})$, $k\in\mathds{N}$. One can see that $1\leq\beta_{k+1}\leq\beta_{k}$ as well as $\|x^{k}-w^{0}\|\geq 1$, for all $k\in\mathds{N}$, so that the sequence $(x^{k})_{k\in\mathds{N}}$ is well-defined.
Now, consider $\lambda\in(0,1]$ and set $w\coloneqq(\lambda,0)\in\operatorname{conv}(M)$. It so happens that $\|x^{k+1}-w\|\leq\|x^{k}-w\|$, for all odd $k$. Indeed,
$\beta_{k+1}=\sqrt{\alpha_{k}^{2}-2\alpha_{k}+\beta_{k}^{2}}\leq\sqrt{\alpha_{k}^{2}-2\alpha_{k}\lambda+\beta_{k}^{2}}\eqqcolon\gamma_{k},$
and $\gamma_{k}$ is such that $z^{k}\coloneqq(0,\gamma_{k})$ belongs to the circumference centered in $w$ and passing through $x^{k}$. In other words, $z^{k}$ complies with $\|z^{k}-w\|=\|x^{k}-w\|$. Thus,
$\|x^{k+1}-w\|=\sqrt{\beta_{k+1}^{2}+\lambda^{2}}\leq\sqrt{\gamma_{k}^{2}+\lambda^{2}}=\|x^{k}-w\|.$
On the other hand, for $w^{\ell}=(1/2^{\ell},0)\in{M}$ with $\ell\geq 0$, consider $N(w^{\ell})=2\ell$. For all even $k\geq N(w^{\ell})$, we have $\alpha_{k}=0$, $\alpha_{k+1}\leq 1/2^{\ell}$ and $\beta_{k+1}=\beta_{k}$. So, $\|x^{k+1}-w^{\ell}\|\leq\|x^{k}-w^{\ell}\|$ and the sequence is Fejér* monotone w.r.t. $M$.
Note that $\operatorname{conv}(M)=(0,1]\times\{0\}$ and also it holds that $\bar{w}=(0,0)\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$. We show now that $(x^{k})_{k\in\mathds{N}}$ is not Fejér* monotone w.r.t. $\{(0,0)\}$. Indeed, for $k=2\ell$, $\ell\in\mathds{N}$, we have $x^{2\ell}=(0,\beta_{2\ell})$ and $x^{{2\ell}+1}=(1/2^{\ell},\beta_{{2\ell}+1})=(1/2^{\ell},\beta_{2\ell})$. Thus,
$\|x^{{2\ell}+1}-\bar{w}\|^{2}=\dfrac{1}{2^{2\ell}}+\beta_{2\ell}^{2}>\beta_{2\ell}^{2}=\|x^{2\ell}-\bar{w}\|^{2}.$
Finally, just for the sake of completion, we show that sequence $(x^{k})_{k\in\mathds{N}}$ converges to $x^{*}\coloneqq(0,\sqrt{4/3})$. In fact, for $k=2\ell$, $\ell\in\mathds{N}$, we have from Example 1 that
$\beta_{2\ell+2}^{2}$ $=\alpha_{2\ell+1}^{2}-2\alpha_{2\ell+1}+\beta_{2\ell+1}^{2}=\dfrac{1}{2^{2\ell}}-2\dfrac{1}{2^{\ell}}+\beta_{2\ell}^{2}$ (4)
$=\displaystyle\sum\limits_{j=0}^{\ell}\dfrac{1}{2^{2j}}-2\sum\limits_{j=0}^{\ell}\dfrac{1}{2^{j}}+\beta_{0}^{2}=\displaystyle\sum_{j=0}^{\ell}\dfrac{1}{2^{2j}}\stackrel{{\scriptstyle\ell\to+\infty}}{{\to}}\dfrac{4}{3}.$ (5)
Clearly, we have $\left\lVert x^{2\ell+1}-x^{2\ell}\right\rVert\to 0$ when $\ell\to+\infty$, so $x^{k}\to x^{*}$, as desired. ∎
Remark 1
One might argue that the set $M$ in Example 1 is somehow special, for $M$ has empty interior. However, it is possible to prove the same result for an extended example where the underlying set has nonempty interior, say, for instance, $M\coloneqq\{(w^{\prime},w^{\prime\prime})\in\mathds{R}^{2}\mid 0<w^{\prime}\leq 1,-1<w^{\prime\prime}\leq 0\}$.
In the next proposition the closure of the convex hull of the underlying set plays a role. Even though a Fejér* monotone sequence w.r.t. $M\subset\mathcal{H}$ is not Fejér* monotone w.r.t. $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, one can characterize Fejér* monotonicity in terms of the convergence of scalar sequences and inner product sequences to the closure of the convex hull. This result can be useful for the analysis regarding weak convergence of optimization algorithms that generate Fejér* monotone sequences.
Proposition 2
Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be a Fejér* monotone w.r.t. a nonempty set $M$ in $\mathcal{H}$. Then,
  1. (i)
    for every $\bar{x}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$ the scalar sequence $(\left\lVert x^{k}-\bar{x}\right\rVert)_{k\in\mathds{N}}$ converges;
  2. (ii)
    for every $\bar{x}_{1},\bar{x}_{2}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, it holds that the sequence $\left(\left\langle\bar{x}_{1}-\bar{x}_{2},x^{k}\right\rangle\right)_{k\in\mathds{N}}$ converges.
Proof
Regarding item (i), recall that Proposition 1(iii) implies that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $\operatorname{conv}(M)$. Thus, for any $w\in\operatorname{conv}(M)$, Proposition 1(ii) yields that the scalar sequence $(\left\lVert x^{k}-w\right\rVert)_{k\in\mathds{N}}$ converges.
Now, take $\bar{x}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, and consider a sequence $(w^{\ell})_{k\in\mathds{N}}\subset\operatorname{conv}(M)$ such that $w^{\ell}\rightarrow\bar{x}$. Using the triangle inequality, we have, for all $\ell\in\mathds{N}$,
$-\left\|w^{\ell}-\bar{x}\right\|\leq\left\|x^{k}-\bar{x}\right\|-\left\|x^{k}-w^{\ell}\right\|\leq\left\|w^{\ell}-\bar{x}\right\|.$ (6)
[…]
$-\left\|w^{\ell}-\bar{x}\right\|$ $\leq\liminf_{k}\left\|x^{k}-\bar{x}\right\|-\lim_{k}\left\|x^{k}-w^{\ell}\right\|$ (7)
$\leq\limsup_{k}\left\lVert x^{k}-\bar{x}\right\rVert-\lim_{k}\|x^{k}-w^{\ell}\|\leq\left\|w^{\ell}-\bar{x}\right\|,$ (8)
because, as aforementioned, item Proposition 1(ii) says that, for every $\ell\in\mathds{N}$, $\lim_{k}\left\|w^{\ell}-x^{k}\right\|$ exists. Now, as $\ell$ goes to infinity, we have $\left\|w^{\ell}-\bar{x}\right\|\to 0$, and we conclude that
$\liminf_{k}\left\|x^{k}-\bar{x}\right\|=\limsup_{k}\left\|x^{k}-\bar{x}\right\|=\lim_{k}\left\|x^{k}-\bar{x}\right\|,$
so that we have the desired result.
As for item (ii), let $\bar{x}_{1},\bar{x}_{2}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$. Then,
$\left\langle\bar{x}_{1}-\bar{x}_{2},x^{k}\right\rangle$ $=\left\langle\bar{x}_{1}-\bar{x}_{2},x^{k}-\bar{x}_{1}\right\rangle+\left\langle\bar{x}_{1}-\bar{x}_{2},\bar{x}_{1}\right\rangle$ (9)
$=\frac{1}{2}\left(\left\|x^{k}-\bar{x}_{2}\right\|^{2}-\left\|\bar{x}_{1}-\bar{x}_{2}\right\|^{2}-\left\|x^{k}-\bar{x}_{1}\right\|^{2}\right)+\left\langle\bar{x}_{1}-\bar{x}_{2},\bar{x}_{1}\right\rangle.$ (10)
Now, item (i) implies that the sequences $(\left\|x^{k}-\bar{x}_{1}\right\|)_{k\in\mathds{N}}$ and $(\left\|x^{k}-\bar{x}_{2}\right\|)_{k\in\mathds{N}}$ converge. Therefore, the sequence $\left(\left\langle\bar{x}_{1}-\bar{x}_{2},x^{k}\right\rangle\right)_{k\in\mathds{N}}$ converges. ∎
4 Fejér* and quasi-Fejér sequences
In this section we discuss Fejér* monotonicity relation with quasi-Fejér monotone sequences (as addressed in Combettes:2001a ). First, we establish the following notation: $\ell_{+}$ is the set of real sequences with nonnegative entries, and $\ell^{1}$ is the set of real sequences with finite $\ell^{1}$-norm.
Definition 2 (quasi-Fejér monotonicity (Combettes:2001a, , Def. 1.1))
Let $M\subset\mathcal{H}$ be a nonempty set. We say that a sequence $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$, with respect to $M$, is
  1. (i)
    quasi-Fejér of Type I if there exists a sequence $(\varepsilon_{k})_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$ such that, for all $x\in M$ and $k\in\mathds{N}$, it holds that
    $\left\|x^{k+1}-x\right\|\leq\left\|x^{k}-x\right\|+\varepsilon_{k};$
  2. (ii)
    quasi-Fejér of Type II if there exists a sequence $(\varepsilon_{k})_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$ such that, for all $x\in M$ and $k\in\mathds{N}$, it holds that
    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}+\varepsilon_{k};$
  3. (iii)
    quasi-Fejér of Type III if for all $x\in M$ there exists a sequence $(\varepsilon_{k})_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$ such that, for all $k\in\mathds{N}$, it holds that
    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}+\varepsilon_{k}.$
Sure enough, Fejér sequences are quasi-Fejér of Type I sequences. Also, one can clearly see that Type II implies Type III. Moreover, it was established in (Combettes:2001a, , Prop. 3.2) that if $M$ is bounded, then Type I implies Type II.
Quasi-Fejér of Type III sequences are the most general ones. However, due to this generality, they are hardly used for the analysis of optimization algorithms. In fact, the most common type of quasi-Fejér sequences used in the literature are quasi-Fejér of Type II sequences; see, for instance, Alber:1998 ; Iusem:1994 .
The next theorem shows that Fejér* monotonicity sequences are also quasi-Fejér of Type III.
Theorem 4.1
Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be a Fejér* monotone sequence w.r.t. a nonempty set $M$ in $\mathcal{H}$. Then, $(x^{k})_{k\in\mathds{N}}$ is a quasi-Fejér of Type III sequence w.r.t. $M$.
Proof
Suppose that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $M$. Then, in view of Definition 1, for any $x\in M$, there exists $N(x)$ such that, for all $k\geq N(x)$, it holds that
$\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}.$
Let us now define the sequence $(\varepsilon_{k})_{k\in\mathds{N}}$. For $k\geq N(x)$, we set $\varepsilon_{k}\coloneqq 0$. On the other hand, for $k<N(x)$, we might have
$\left\|x^{k+1}-x\right\|>\left\|x^{k}-x\right\|,$
so we define
$\varepsilon_{k}\coloneqq\begin{cases*}\left\|x^{k+1}-x\right\|^{2}-\left\|x^{k}-x\right\|^{2},&\text{if $\left\|x^{k+1}-x\right\|>\left\|x^{k}-x\right\|$},\\ 0,&\text{otherwise}.\end{cases*}$
Thus, it holds true that sequence $(\varepsilon_{k})_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$ and therefore sequence $(x^{k})_{k\in\mathds{N}}$ is quasi-Fejér of Type III w.r.t. $M$. ∎
We note that the concepts of Fejér* and quasi-Fejér of Type III sequences clearly are not equivalent, with the latter being more general than the former.
Actually, if $M$ is a singleton, then any Fejér* monotone sequence is quasi-Fejér of Type II, using the same arguments used in the proof of Theorem 4.1. Of course, in this case one can easily build quasi-Fejér sequences of Type II that are not Fejér*. Nevertheless, if a sequence is quasi-Fejér of Type II w.r.t. a general $M$, then it is Type II with respect to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$ (see (Combettes:2001a, , Prop. 3.4)). Hence, in view of Example 1, we have a Fejér* sequence that it is not quasi-Fejér of Type II monotone. Note that we can construct a non-singleton set $M$, for instance, $M\coloneqq[-1,0]\times\{0\}\subset\mathds{R}^{2}$, such that the same sequence $(x^{k})_{k\in\mathds{N}}$ of Example 1 is quasi-Fejér of Type II, but not Fejér* monotone, w.r.t. $M$. More than not being equivalent, we have just seen that the concepts of Fejér* and quasi-Fejér of Type II are not contained in each other, that is, those ideas are independent.
The previous theorem reveals a particular quasi-Fejér of type III sequence that is related to optimization algorithms, through Fejér* monotonicity Behling:2024c . Thus, we end our discussion by presenting some characterizations of Fejér* monotonicity in terms of the weak convergence of the underlying sequence. These statements are inspired by Combettes:2001a and could be derived from the results therein, since Fejér* monotonicity implies quasi-Fejér of type III monotonicity. However, we present them here, for the sake of completeness. Moreover, the provided proofs are simpler and more clear for the reader, in the context of Fejér* monotonicity.
From now on, for a sequence $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$, we denote by $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$ the set of its weak cluster points and by $\mathfrak{S}\left(x^{k}\right)_{k\in\mathds{N}}$ the set of its strong cluster points. First, we give portrayals of Fejér* monotonicity in terms of the weak convergence of the sequence.
Proposition 3
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to $M\subset\mathcal{H}$. Then,
  1. (i)
    $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\neq\emptyset$.
  2. (ii)
    for any $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$, there exists an $\alpha\in\mathds{R}$ such that
    $M\subset\left\{y\in\mathcal{H}\mid\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha\right\};$
  3. (iii)
    if $\operatorname{cl}\left({\operatorname{aff}(M)}\right)=\mathcal{H}$, then $\left(x^{k}\right)_{k\in\mathds{N}}$ converges weakly;
  4. (iv)
    if $x^{k}\rightharpoonup\bar{x}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, then the scalar sequence $\left(\left\|x^{k}-y\right\|\right)_{k\in\mathds{N}}$ converges for every $y\in\mathcal{H}$.
Proof
Item (i): the statement is valid since, due to Proposition 1(i), sequence $\left(x^{k}\right)_{k\in\mathds{N}}$ is bounded.
Item (ii): Take any $y\in M$. For all $k\in\mathds{N}$, we get
$\left\|x^{k}-y\right\|^{2}-\|y\|^{2}=\left\|x^{k}\right\|^{2}-2\left\langle y,x^{k}\right\rangle.$
Moreover, it follows from Proposition 1(ii) that $\lim\left(\left\|x^{k}-y\right\|^{2}-\|y\|^{2}\right)$ is well-defined.
Now, given two points $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$, such that $x^{j_{k}}\rightharpoonup w^{\prime}$ and $x^{\ell_{k}}\rightharpoonup w^{\prime\prime}$, and considering Section 4 we have
$\lim\left\|x^{j_{k}}\right\|^{2}-2\left\langle y,w^{\prime}\right\rangle=\lim\left\|x^{\ell_{k}}\right\|^{2}-2\left\langle y,w^{\prime\prime}\right\rangle,$
and therefore
$\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\frac{1}{2}\left(\lim\left\|x^{j_{k}}\right\|^{2}-\lim\left\|x^{\ell_{k}}\right\|^{2}\right).$
Setting $\alpha$ as the right-hand side of the above equation yields that
$M\subset\left\{y\in\mathcal{H}\mid\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha\right\},$
as required.
Item (iii): In view of item (ii), if $\operatorname{cl}\left({\operatorname{aff}(M)}\right)=\mathcal{H}$ then, for all $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$, there exists an $\alpha\in\mathds{R}$ such that for all $y\in\mathcal{H}$ it holds that $\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha$. Consequently, $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$ reduces to a singleton. Due to Proposition 2(i), $\left(x^{k}\right)_{k\in\mathds{N}}$ lies in a weakly compact set and therefore converges weakly.
Item (iv): Consider any $y\in\mathcal{H}$. Then, for all $k\in\mathds{N}$, we have
$\left\|x^{k}-y\right\|^{2}=\left\|x^{k}-\bar{x}\right\|^{2}+2\left\langle x^{k}-\bar{x},\bar{x}-y\right\rangle+\|\bar{x}-y\|^{2}.$
Using the hypothesis and Proposition 2 on the right-hand side of the above equation, we conclude the convergence of $\left(\left\|x^{k}-y\right\|\right)_{k\in\mathds{N}}$.
Theorem 4.2
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to a nonempty set $M\subset\mathcal{H}$. Then, $\left(x^{k}\right)_{k\in\mathds{N}}$ converges weakly to a point in $M$ if, and only if, $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\subset M$.
Proof
It holds trivially that Fejér* monotonicity together with weakly convergence implies that $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\subset M$. On the other hand, suppose that $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\subset M$ and take $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$. Thus, $w^{\prime},w^{\prime\prime}\in M$. Proposition 3(ii) affirms that $\left\langle w^{\prime},w^{\prime}-w^{\prime\prime}\right\rangle=\left\langle w^{\prime\prime},w^{\prime}-w^{\prime\prime}\right\rangle$ and hence, $w^{\prime}=w^{\prime\prime}$. The boundedness of $\left(x^{k}\right)_{k\in\mathds{N}}$, guaranteed by Proposition 1, completes the proof. ∎
We conclude this section by presenting a pair of results that characterizes the strong convergence of Fejér* monotone sequences.
Proposition 4
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to a nonempty set $M\subset\mathcal{H}$. Then,
  1. (i)
    for all $w^{\prime},w^{\prime\prime}\in\mathfrak{S}\left(x^{k}\right)_{k\in\mathds{N}}$, it holds that
    $M\subset\left\{y\in\mathcal{H}\mid\left\langle y-\frac{w^{\prime}+w^{\prime\prime}}{2},w^{\prime}-w^{\prime\prime}\right\rangle=0\right\};$
  2. (ii)
    if $\operatorname{cl}\left({\operatorname{aff}(M)}\right)=\mathcal{H}$, then $\mathfrak{S}\left(x^{k}\right)_{k\in\mathds{N}}$ contains at most one point;
  3. (iii)
    the sequence converges strongly if there exist $x\in M,\left(\varepsilon_{k}\right)_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$, and $\rho\in(0,+\infty)$ such that for all $k\in\mathds{N}$ it holds that
    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}-\rho\left\|x^{k+1}-x^{k}\right\|+\varepsilon_{k}.$
Proof
It follows from Theorem 4.1 together with Proposition 3.7 in Combettes:2001a . ∎
Theorem 4.3
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to a nonempty set $M\subset\mathcal{H}$. Then, the following statements are equivalent:
  1. (i)
    $\left(x^{k}\right)_{k\in\mathds{N}}$ converges strongly to a point in $M$;
  2. (ii)
    $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\subset M$ and $\mathfrak{S}\left(x^{k}\right)_{k\in\mathds{N}}\neq\varnothing$;
  3. (iii)
    $\mathfrak{S}\left(x^{k}\right)_{k\in\mathds{N}}\cap M\neq\varnothing$.
Proof
This is a direct result of combining Theorem 4.1 and Theorem 3.11 in Combettes:2001a . ∎
We emphasize that new characterizations of Fejér* monotonicity could be revealed with further analysis, suggesting a direction for future research.
5 Concluding remarks
In this work we further expand the understanding of Fejér* monotonicity. We provide a characterization of Fejér* monotonicity in terms of the convergence of scalar sequences and inner product sequences to the closure of the convex hull of the underlying set. We provide an insightful example that shows that Fejér* monotonicity not always can be extended to the closure of the convex hull of the underlying set. It was also established connections between Fejér* monotonicity and quasi-Fejér monotonicity. The results presented in this paper could shed a light on the analysis of the (weak) convergence of algorithms that comply with Fejér* monotonicity.
Acknowledgements.
RB was partially supported by Brazilian agency Fundação de Amparo à Pesquisa do Estado do Rio de Janeiro (Grant E-26/201.345/2021); YBC was partially supported by the USA agency National Science Foundation (Grant DMS-2307328); AAR was partially supported by Brazilian agency Conselho Nacional de Desenvolvimento Científico e Tecnológico – CNPq (Grant 307987/2023-0); LRS was also partially funded by CNPq (Grant 310571/2023-5).

### v2

@@
-
+Fejér* monotonicity in optimization algorithms
 Abstract
-Fejér monotonicity is a well-established property commonly observed in sequences generated by optimization algorithms.
-In this paper, we introduce an extension of this property, called Fejér* monotonicity, which was initially proposed in [SIAM J.
+Fejér monotonicity is a well-established property often observed in sequences generated by optimization algorithms.
+In this paper, we study an extension of this property, called Fejér* monotonicity, which was initially proposed in [SIAM J.
 Optim., 34(3), 2535–2556 (2024)].
-We discuss and build upon the concept by exploring its behavior within Hilbert spaces, presenting an illustrative example and insightful results regarding weak and strong convergence.
-We also compare Fejér* monotonicity with other weak notions of Fejér-like monotonicity, to better establish the role of Fejér* monotonicity in optimization algorithms.
+We discuss and explore its behavior within Hilbert spaces as a tool for optimization algorithms.
+Additionally, we investigate weak and strong convergence properties of this novel concept.
+Through illustrative examples and insightful results, we contrast Fejér* with weaker notions of quasi-Fejér-type monotonicity.
 keywords:
-Fejér monotonicity convex analysis optimization algorithms
-MSC:
-49M27 65K05 65B99 90C25
-journal: JOTA
+Fejér* monotonicity, Fejér monotonicity, convex analysis, optimization algorithms
+pacs:
+MSC (2020)]49M27, 65K05, 90C25
 1 Introduction
 A useful concept for analyzing the convergence of optimization algorithms is the notion of Fejér monotonicity.
-A sequence $(x^{k})_{k\in\mathds{N}}$ defined in a Hilbert space $\mathcal{H}$ is said to be Fejér monotone with respect to a nonempty set $M\subset\mathcal{H}$ if for all $x\in M$ and all $k\in\mathds{N}$ we have $\lVert x^{k+1}-x\rVert\leq\lVert x^{k}-x\rVert$.
-It is well-known that if $(x^{k})_{k\in\mathds{N}}$ is Fejér monotone with respect to $M$, then $(x^{k})_{k\in\mathds{N}}$ is bounded, and if all its weak cluster points belong to $M$, then the entire sequence $(x^{k})_{k\in\mathds{N}}$ converges weakly to some point lying in $M$; see, for instance, (Bauschke:1996, , Theorem 2.16), and Combettes:2009 for further details and historical notes.
-The notion of Fejér monotonicity is named after Hungarian mathematician Lipót Fejér (1880-1959) who introduced this notion in Fejer:1922 and further developed in Fejer:1937 and later in Agmon:1954 ; Motzkin:1954 .
+We denote by $\mathds{N}$ the set of all nonnegative integers.
+A sequence $(x^{k})_{k\in\mathds{N}}$ defined in a real Hilbert space $\mathcal{H}$ is said to be Fejér monotone with respect to a nonempty set $M\subset\mathcal{H}$ if for all $x\in M$ and all $k\in\mathds{N}$ we have $\lVert x^{k+1}-x\rVert\leq\lVert x^{k}-x\rVert$.
+It is well-known that if $(x^{k})_{k\in\mathds{N}}$ is Fejér monotone with respect to $M$, then $(x^{k})_{k\in\mathds{N}}$ is bounded, and if all its weak cluster points belong to $M$, then the entire sequence $(x^{k})_{k\in\mathds{N}}$ converges weakly to some point lying in $M$; see, for instance, [Bauschke:1996, Thm 2.16], [Bauschke:2017a, Prop.
+5.4 and Thm.
+5.5, p.
+92], and [Combettes:2009] for further details and historical notes.
+The notion of Fejér monotonicity is named after Hungarian mathematician Lipót Fejér (1880-1959) who introduced this notion in [Fejer:1922] and further developed it in [Fejer:1937] and later in [Agmon:1954, Motzkin:1954].
 Fejér was the force behind the highly successful Hungarian school of analysis.
@@
 Weaker notions of Fejér monotonicity have been considered in the literature.
-For instance, the concept of quasi-Fejér monotonicity, first presented in Ermolev:1969 and further elaborated in several works such as Iusem:1994 ; Combettes:2001a .
+For instance, the concept of quasi-Fejér monotonicity, first presented in [Ermolev:1969] and further elaborated in several works such as [Iusem:1994, Combettes:2001a].
 In this work, we explore a slightly different notion, called Fejér* monotonicity.
-It was first introduced inBehling:2024c as a key ingredient to show that the Circumcentered-Reflection Method (CRM) Behling:2024b , proposed by these authors, solves the Convex Feasibility Problem (CFP).
+It was first introduced in [Behling:2024c] as a key ingredient to show that the Circumcentered-Reflection Method (CRM) [Behling:2024b], proposed by the authors of [Behling:2024c], solves the Convex Feasibility Problem (CFP).
 The main difference with the usual Fejér monotonicity notion relies on the fact that now the nonincreasing distance property holds for the tail of the sequence, starting at some index which depends on the considered point $x\in M$.
+In this work we further explore this novel concept.
+Indeed, the main contributions of this paper are twofold: (i) We present a detailed discussion on the properties of Fejér* monotonicity, including the ones that can be used as a tool to analyze convergence of sequences generated by optimization algorithms as well as the relationship of Fejér* with other weaker notions of Fejér monotonicity; (ii) We provide an illustrative example that highlights the differences between Fejér* and quasi-Fejér monotonicity.
+This example is key to clarify the connections and differences of these concepts.
 This paper is organized as follows.
-We start in Section 2 with a brief discussion on the role of Fejér monotonicity in optimization algorithms.
-In Section 3 we present some basic properties of Fejér* monotonicity, and we make a discussion, upon an illustrative example on properties of Fejér* monotonicity.
-In Section 4 we compare Fejér* sequences and quasi-Fejér sequences.
-Finally, in Section 5 we present some concluding remarks.
+We start in Section˜2 with a brief discussion on the role of Fejér monotonicity in optimization algorithms.
+In Section˜3 we present some basic properties of Fejér* monotonicity, and we discuss, upon illustrative examples, some properties of Fejér* monotonicity sequences.
+In Section˜4 we compare Fejér* sequences and quasi-Fejér sequences.
+Finally, in Section˜5 we present some concluding remarks.
 2 Fejér* monotonicity in optimization algorithms
 First, let us formally introduce the notion of Fejér* monotonicity.
-Definition 1 (Fejér* monotonicity)
+Definition 2.1 (Fejér* monotonicity).
 Let $M\subset\mathcal{H}$ be a nonempty set and consider a sequence $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$.
 We say that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $M$ if for any point $x\in M$, there exists $N(x)\in\mathds{N}$ such that, for all $k\geq N(x)$,
-$\left\lVert x^{k+1}-x\right\rVert\leq\left\lVert x^{k}-x\right\rVert.$
-We address now the role of Fejér and Fejér monotonicity in connection with fixed-point and feasibility problems.
-Consider an operator $T:\mathds{R}^{n}\to\mathds{R}^{n}$.
+$\|x^{k+1}-x\|\leq\|x^{k}-x\|.$ (2)
+Note that Fejér monotonicity implies Fejér* monotonicity, but the converse is not true in general.
+We address now the role of Fejér and Fejér* monotonicity in connection with fixed-point and feasibility problems.
+Consider an operator $T:\mathcal{H}\to\mathcal{H}$.
 We remind that $T$ is a contraction if there exists $\tau\in(0,1)$ such that
-$\left\lVert T(x)-T(y)\right\rVert\leq\tau\left\lVert x-y\right\rVert$
-for all $x,y\in\mathds{R}^{n}$.
-It is well known, and elementary, that a contraction has a unique fixed point (i.e., a point $\bar{x}\in\mathds{R}^{n}$ such that $T(\bar{x})=\bar{x}$) and that the sequence obtained by iterating the operator (i.e., the sequence $(x^{k})_{k\in\mathds{N}}\subset\mathds{R}^{n}$ defined as $x^{k+1}=T(x^{k})$ starting at any $x^{0}\in\mathds{R}^{n}$), converges to such unique fixed point.
+$\left\lVert T(x)-T(y)\right\rVert\leq\tau\left\lVert x-y\right\rVert,$ (3)
+for all $x,y\in\mathcal{H}$.
+It is well known that a contraction $T$ has a unique fixed point (i.e., a point $\bar{x}\in\mathcal{H}$ such that $T(\bar{x})=\bar{x}$) and that the sequence obtained by iterating the operator $T$ (i.e., the sequence $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ defined as $x^{k+1}=T(x^{k})$ starting at any $x^{0}\in\mathcal{H}$), converges to such a unique fixed point; for further details on contraction mappings, see [Kirk:2001].
 The situation changes when we consider notions which are weaker than the contraction property. The most basic one is that of nonexpansiveness. An operator $T$ is nonexpansive if
-$\left\lVert T(x)-T(y)\right\rVert\leq\left\lVert x-y\right\rVert$
-for all $x,y\in\mathds{R}^{n}$.
-The set of fixed points of a nonexpansive operator $T$ may be empty, or infinite, and the behavior of the sequence $(x^{k})_{k\in\mathds{N}}$ obtained by iterating $T$ may fail to converge, but given any fixed point $\bar{x}$ of $T$, it follows immediately from Section 2 that $\left\lVert x^{k+1}-\bar{x}\right\rVert\leq\left\lVert x^{k}-\bar{x}\right\rVert$, i.e., the sequence is Fejér monotone to the set $M$ of fixed points of $T$, and thus it enjoys the basic properties of Fejér monotone sequences, namely that it is bounded, and that if some of its cluster points belongs to $M$, then the sequence $(x^{k})_{k\in\mathds{N}}$ converges to it; see, for instance, Bauschke:2017a for more on nonexpansiveness and Fejér monotonicity.
+$\left\lVert T(x)-T(y)\right\rVert\leq\left\lVert x-y\right\rVert,$ (4)
+for all $x,y\in\mathcal{H}$.
+The set of fixed points of a nonexpansive operator $T$ may be empty, or infinite, and the behavior of the sequence $(x^{k})_{k\in\mathds{N}}$ obtained by iterating $T$ may fail to converge, but given any fixed point $\bar{x}$ of $T$, it follows immediately from Eq.˜4 that $\left\lVert x^{k+1}-\bar{x}\right\rVert\leq\left\lVert x^{k}-\bar{x}\right\rVert$, i.e., the sequence is Fejér monotone with respect to the set $M$ of fixed points of $T$, and thus it enjoys the basic properties of Fejér monotone sequences, namely that it is bounded, and that if all its cluster points belong to $M$, then the sequence $(x^{k})_{k\in\mathds{N}}$ converges to one of them; see, for instance, [Bauschke:2017a] for more on nonexpansiveness and Fejér monotonicity.
 Among the nonexpansive operators, projections onto convex sets are quite significant.
-We remind that, given a closed and convex set $C\subset\mathds{R}^{n}$, the projection onto $C$, $P_{C}:\mathds{R}^{n}\to C$, is defined as $P_{C}(x)\coloneqq\operatorname*{arg\,min}_{y\in C}\left\lVert x-y\right\rVert$.
-Convex combinations of projections (i.e., operators of the form $\bar{P}=\sum_{i+1}^{m}\alpha_{i}P_{C_{i}}$ with $\alpha_{i}\in[0,1],\sum_{i=1}^{m}\alpha_{i}=1$, and $C_{i}\subset\mathds{R}^{n}$ convex) are nonexpansive, and the same holds for compositions of projections, i.e., operator of the form $\widehat{P}=P_{C_{m}}\circ\dots\circ P_{C_{1}}$.
-The set of fixed points of a projection $P_{C}$ is $C$, and the sets of fixed points of $\bar{P}$ and $\widehat{P}$, as defined above are equal to the set $C:=\cap_{i=1}^{m}C_{i}$.
-Iterating the operators $\bar{P}$ and $\widehat{P}$, one obtains the so-called Simultaneous Projections and Sequential Projections algorithms, which aim to solve the (CFP), consisting of finding a point in the intersection $C$ of closed and convex sets $C_{1},\dots,C_{m}$.
-See Bauschke:1996 for a detailed discussion on these methods, which can be traced back to the seminal works of Cimmino Cimmino:1938 and Kaczmark Kaczmarz:1937 .
-In view of the discussion above, the sequences generated by these two methods are Fejér monotone to $C$.
+We remind that, given a nonempty, convex and closed set $C\subset\mathcal{H}$, the projection onto $C$, $P_{C}:\mathcal{H}\to C$, is defined as $P_{C}(x)\coloneqq\operatorname*{arg\,min}_{y\in C}\left\lVert x-y\right\rVert$.
+Convex combinations of projections (i.e., operators of the form $\bar{P}=\sum_{i=1}^{m}\alpha_{i}P_{C_{i}}$ with $\alpha_{i}\in[0,1],\sum_{i=1}^{m}\alpha_{i}=1$, and $C_{i}\subset\mathcal{H}$ convex) are nonexpansive, and the same holds for compositions of projections, i.e., operator of the form $\widehat{P}=P_{C_{m}}\circ\dots\circ P_{C_{1}}$.
+The set of fixed points of a projection $P_{C}$ is $C$, and the sets of fixed points of $\bar{P}$ and $\widehat{P}$, as defined above, are equal to the set $C:=\cap_{i=1}^{m}C_{i}$ whenever $C\neq\emptyset$.
+Iterating the operators $\bar{P}$ and $\widehat{P}$, one obtains the so-called Simultaneous Projections and Sequential Projections algorithms, which aim to solve the CFP, consisting of finding a point in the intersection $C$ of nonempty, convex and closed sets $C_{1},\dots,C_{m}$.
+See the work of Bauschke:1996 for a detailed discussion on these methods, which can be traced back to the seminal works of Cimmino:1938 and Kaczmarz:1937.
+In view of the discussion above, the sequences generated by these two methods are Fejér monotone with respect to $C$.
 In order to show that the sequences generated by these algorithms do converge to a point in $C$, one needs a bit more than nonexpansiveness.
@@
     firm nonexpansiveness, meaning that
-    $\left\lVert T(x)-T(y)\right\rVert^{2}\leq\left\lVert x-y\right\rVert^{2}-\left\lVert(T(x)-T(y))-(x-y)\right\rVert^{2}$
-    for all $x,y\in\mathds{R}^{n}$; and
+    $\left\lVert T(x)-T(y)\right\rVert^{2}\leq\left\lVert x-y\right\rVert^{2}-\left\lVert(T(x)-T(y))-(x-y)\right\rVert^{2}$ (5)
+    for all $x,y\in\mathcal{H}$; and
   2. (ii)
     nonexpansiveness plus, meaning that
-    $\left\lVert T(x)-T(y)\right\rVert\leq\left\lVert x-y\right\rVert$
-    for all $x,y\in\mathds{R}^{n}$, and $T(x)-T(y)=x-y$ whenever equality holds in (ii).
-Projections onto convex sets, as well as the above defined operators $\bar{P}$ and $\widehat{P}$, are firmly nonexpansive.
-Convex combinations of firmly nonexpansive operators may fail to be firmly nonexpansive, but they are nonexpansive plus (see Arefidamghani:2023 ).
-With the help of the Fejér convergence properties and the notion of firm nonexpansiveness, it is easy to prove that the Simultaneous and the Successive Projection algorithms do converge to a point in $C$, whenever $C$ is nonempty (see Bauschke:1996 ).
+    $\left\lVert T(x)-T(y)\right\rVert\leq\left\lVert x-y\right\rVert$ (6)
+    for all $x,y\in\mathcal{H}$, and T(x)T(y)=xyT(x)-T(y)=x-y whenever equality holds in Eq.˜6.
+Projections onto nonempty, convex and closed sets, as well as the above defined operator $\bar{P}$, are firmly nonexpansive.
+With the help of Fejér convergence properties and the notion of firm nonexpansiveness, it is easy to prove that the Simultaneous and the Successive Projection algorithms do converge to a point in $C$, whenever $C$ is nonempty (see [Bauschke:1996]).
 One way to accelerate these methods for CFP is to approximate the set $C$ from the inside, i.e., to replace at iteration $k$ the set $C$ by a set $C^{k}$ such that
-$C^{k}\subset C^{k+1}\subset\dots\subset C,$
-for all $k$. For instance, if $C\coloneqq\{x\in\mathds{R}^{n}\mid g(x)\leq 0\}$ for some convex $g:\mathds{R}^{n}\to\mathds{R}$, then $C^{k}$ could be chosen as
-$C^{k}\coloneqq\{x\in\mathds{R}^{n}\mid g(x)+\epsilon_{k}\leq 0\},$
-where $(\epsilon_{k})_{k\in\mathds{N}}$ is a sequence of positive real numbers decreasing to 00.
-Then one could consider the sequence $(x^{k})_{k\in\mathds{N}}$ given by $x^{k+1}=T_{k}(x^{k})$, where $T_{k}$ is the projection onto a convex set which contains $C^{k}$.
-In such a case, the nonexpansiveness of $T^{k}$ gives $\left\lVert x^{k+1}-y\right\rVert\leq\left\lVert x^{k}-y\right\rVert$ but only for points $y\in C^{k}$, not in the whole set $C$.
-However, if we take $C^{k}$ as in Section 2, and $C$ has nonempty interior, any point $y\in$ int$(C)$ (i.e., such that $g(y)<0$), will belong to $C^{k}$ for all $k$ bigger than $k(y)$, defined as the first integer such that $\epsilon_{k(y)}<-g(y)$, due to Section 2, but this is exactly the same as saying that $(x^{k})_{k\in\mathds{N}}$ is Fejér monotone to int$(C)$.
-Hence, Fejér* monotonicity is precisely what is needed for the analysis of algorithms which use approximation schemes as in Section 2.
-Among them, we mention the algorithms in Iusem:1986a ; Iusem:1986 for CFP.
-These methods employ the scheme in Section 2 combined with operators related to $\bar{P}$ and $\widehat{P}$, where the projections onto the sets $C_{i}$, approximated from the inside by sets $C^{k}_{i}$, as in Section 2, are replaced by projections onto hyperplanes separating $x^{k}$ from $C^{k}_{i}$.
-In these cases, the Fejér monotonicity ensures convergence to a point lying possibly in the boundary of $C$ (despite the fact that the sequence is Fejér monotone to int$(C)$).
-A similar situation occurs with the algorithm in Behling:2024c , where the algorithm in Iusem:1986a is enhanced through the addition of a circumcentering step, using the machinery developed during the last few years in Araujo:2022 ; Arefidamghani:2021 ; Arefidamghani:2023 ; Behling:2018 ; Behling:2018a ; Behling:2020 ; Behling:2021b ; Behling:2023 ; Behling:2024 ; Behling:2024b ; Behling:2024c .
-It is worthwhile to note that in the case of Iusem:1986 ; Iusem:1986a ; Behling:2024c , under a Slater hypothesis (i.e., when $\operatorname{int}(C)$ is nonempty), if the $\epsilon_{k}$’s decrease to 00 slowly enough, e.g., so that $\sum_{k=0}^{\infty}\epsilon_{k}=+\infty$, for instance $\epsilon_{k}=1/k$, convergence is indeed finite, i.e., there exists $\bar{k}$ such that $x^{\bar{k}}$ belongs to $C=\bigcap_{i=1}^{m}C_{i}$.
+$C^{k}\subset C^{k+1}\subset\dots\subset C,$ (7)
+for all $k$.
+Then one may consider the sequence $(x^{k})_{k\in\mathds{N}}$ given by $x^{k+1}\coloneqq T_{k}(x^{k})$, where $T_{k}$ is the projection onto a nonempty, convex and closed set which contains $C^{k}$ (but may not be contained or even contain $C$).
+In such a case, the nonexpansiveness of $T_{k}$ gives $\left\lVert x^{k+1}-y\right\rVert\leq\left\lVert x^{k}-y\right\rVert$, but only for points $y\in C^{k}$, not in the whole set $C$.
+For instance, suppose that we have $C\coloneqq\{x\in\mathcal{H}\mid g(x)\leq 0\}$, for some function $g:\mathcal{H}\to\mathds{R}$ that is strictly convex.
+This choice is consistent with the Slater hypothesis (i.e., the existence of $\hat{y}\in\mathcal{H}$, such that $g(\hat{y})<0$) that is needed in this context.
+Let us define
+$C^{k}\coloneqq\{x\in\mathcal{H}\mid g(x)+\epsilon_{k}\leq 0\},$ (8)
+[…]
+Among them, we mention the algorithms in [fukushimaFinitelyConvergentAlgorithm1982, Iusem:1986a, iusemFinitelyConvergentIterative1987, DePierro:1988a], which were proposed for solving CFP.
+These methods employ the scheme in Eq.˜7 combined with operators related to $\bar{P}$ and $\widehat{P}$, where the projections onto the sets $C_{i}$, approximated from the inside by sets $C^{k}_{i}$, as in Eq.˜8, are replaced by projections onto hyperplanes separating $x^{k}$ from $C^{k}_{i}$.
+We stress here that the main Theorem in [fukushimaFinitelyConvergentAlgorithm1982, p.
+1126], Theorem 1 in [iusemFinitelyConvergentIterative1987], and Lemma 1 in [DePierro:1988a] clearly state that only the tail of the sequence generated by these author’s algorithms are Fejér monotone, starting at some index which depends on the considered point.
+Therefore, the sequences are indeed Fejér* monotone.
+A similar situation occurs with the algorithm in [Behling:2024c], where the algorithm in [Iusem:1986a] is enhanced through the addition of a circumcentering step, using the machinery developed during the last few years in [Araujo:2022, Arefidamghani:2021, Arefidamghani:2023, Behling:2018, Behling:2018a, Behling:2020, Behling:2021a, Behling:2023, Behling:2024, Behling:2024b, Behling:2024c].
+It is worthwhile to note that in the case of [Iusem:1986a, Behling:2024c, DePierro:1988a, iusemFinitelyConvergentIterative1987, fukushimaFinitelyConvergentAlgorithm1982], under the Slater hypothesis, if the $\epsilon_{k}$’s decrease to zero slowly enough, e.g., so that $\sum_{k=0}^{\infty}\epsilon_{k}=+\infty$ (take, for instance, $\epsilon_{k}=1/k$), convergence is indeed finite, i.e., there exists $\hat{k}$ such that $x^{\hat{k}}$ belongs to $C=\bigcap_{i=1}^{m}C_{i}$.
+More recently, Kolobov:2021, Kolobov:2022 also make use of the perturbed approximation scheme Eq.˜7 in their algorithms to solve CFPs.
+In particular, Kolobov:2021 also shows that the tail of the sequence generated by their method is Fejér monotone.
+Thus, it is Fejér* monotone.
+Some other properties of Fejér* monotonicity related to optimization algorithms were explored recently by Arakcheev:2025, arakcheevFejerFejerMonotonicity2025, where Fejér* sequences were linked to Opial sequences, and in a survey on Fejér monotonicity by kohlenbachFejerMonotoneSequences2023.
+We refer the reader to these works for further details.
+We highlight that the concept of Fejér* monotonicity is not novel, it has appeared especially when one is dealing with approximations of the feasible set from the inside, and in general is applied in the context of obtaning finite convergence results.
+Our work sheds some light on this concept, exploring its properties and differences with other weaker notions of Fejér monotonicity.
+This study will be carried out in the next sections.
 3 Properties of Fejér* monotonicity
-The first result of this section was already proved in (Behling:2024c, , Prop. 2.5). However, for the sake of completeness, we present it here.
-Proposition 1
-Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be a Fejér* monotone w.r.t. a nonempty set $M$ in $\mathcal{H}$. Then,
+The first result shows some basic properties of Fejér* monotonicity.
+The result was already proved in [Behling:2024c, Prop.
+2.5].
+However, we chose to present the proof here, for the sake of completeness.
+Proposition 3.1.
+Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be Fejér* monotone with respect to a nonempty set $M$ in $\mathcal{H}$. Then,
   1. (i)
@@
   3. (iii)
-    $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $\operatorname{conv}(M)$.
-Proof
+    $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $\operatorname{conv}(M)$.
+Proof.
 For proving (i), take any point $x\in M$.
@@
 Consequently, $(x^{k})_{k\in\mathds{N}}$ is bounded.
-Item (ii) is a direct consequence of Definition 1, since the $N(x)$-tail of sequence $(\left\lVert x^{k}-x\right\rVert)_{k\in\mathds{N}}$ is monotone and bounded, so the sequence converges.
+Item (ii) is a direct consequence of Eq.˜2, since the $N(x)$-tail of sequence $(\left\lVert x^{k}-x\right\rVert)_{k\in\mathds{N}}$ is monotone and bounded, so the sequence converges.
 For item (iii), take any $w\in\operatorname{conv}(M)$.
 Thus, $w$ can be written as $w=\sum_{i=1}^{p}\lambda_{i}w^{i}$, where $w^{i}\in M$, $\lambda_{i}\in[0,1]$ for all $i=1,2,\ldots,p$, $\sum_{i=1}^{p}\lambda_{i}=1$ and $p\in\mathds{N}$.
-Taking into account that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t.
-$M$, for each $i=1,2,\ldots,p$, there exists $N(w^{i})$ such that $\left\lVert x^{k+1}-w^{i}\right\rVert\leq\left\lVert x^{k}-w^{i}\right\rVert$ for all $k\geq N(w^{i})$.
+Taking into account that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $M$, for each $i=1,2,\ldots,p$, there exists $N(w^{i})$ such that $\left\lVert x^{k+1}-w^{i}\right\rVert\leq\left\lVert x^{k}-w^{i}\right\rVert$ for all $k\geq N(w^{i})$.
 Squaring both sides of the last inequality and rearranging the terms, we get
-00 0 $\leq\left\lVert{x^{k}}\right\rVert^{2}-\left\lVert{x^{k+1}}\right\rVert^{2}-2\left\langle x^{k}-x^{k+1},w^{i}\right\rangle.$ (1)
+$0$ $\leq\left\lVert{x^{k}}\right\rVert^{2}-\left\lVert{x^{k+1}}\right\rVert^{2}-2\left\langle x^{k}-x^{k+1},w^{i}\right\rangle.$ (9)
 The above inequality also holds for all $k\geq K\coloneqq\max\{N(w^{1}),N(w^{2}),\ldots,N(w^{p})\}$.
-Multiplying Eq.
-1, for each $i=1,2,\ldots,p$, by the respective $\lambda_{i}$ and adding up, we obtain
-00 0 $\leq\left\lVert{x^{k}}\right\rVert^{2}-\left\lVert{x^{k+1}}\right\rVert^{2}-2\left\langle x^{k}-x^{k+1},\sum_{i=1}^{p}\lambda_{i}w^{i}\right\rangle$ (2)
+Multiplying Eq.˜9, for each $i=1,2,\ldots,p$, by the respective $\lambda_{i}$ and adding up, we obtain
+$0$ $\leq\left\lVert{x^{k}}\right\rVert^{2}-\left\lVert{x^{k+1}}\right\rVert^{2}-2\left\langle x^{k}-x^{k+1},\sum_{i=1}^{p}\lambda_{i}w^{i}\right\rangle$ (10)
 $=\left\lVert{x^{k}}\right\rVert^{2}-\left\lVert{x^{k+1}}\right\rVert^{2}-2\left\langle x^{k}-x^{k+1},w\right\rangle,$ (11)
 and therefore, $\left\lVert x^{k+1}-w\right\rVert\leq\left\lVert x^{k}-w\right\rVert$, for all $k\geq K$.
-So $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t.
-$\operatorname{conv}(M)$, as the proof is complete.
+So $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $\operatorname{conv}(M)$, and the proof is complete.
 ∎
-We remark that a Fejér monotone sequence with respect to a nonempty set $M\subset\mathcal{H}$ is also Fejér monotone with respect to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, the closure of $\operatorname{conv}(M)$; see (Bauschke:2023, , Lem.
-2.1(ii)).
-Furthermore, Proposition 1(iii) guarantees that a Fejér* monotone sequence w.r.t.
-$M$ is as well Fejér* monotone w.r.t.
-$\operatorname{conv}(M)$.
-However, the Fejér* monotonicity cannot always be extended to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, as shown in the next example.
-Example 1
+Remark 3.2.
+We remark that a Fejér monotone sequence with respect to a nonempty set $M\subset\mathcal{H}$ is also Fejér monotone with respect to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, the closure of $\operatorname{conv}(M)$; see Cegielski:2012 and Combettes:2001a, or even Bauschke:2023.
+Furthermore, Proposition˜3.1(iii) guarantees that a Fejér* monotone sequence with respect to $M$ is as well Fejér* monotone with respect to $\operatorname{conv}(M)$.
+However, Fejér* monotonicity cannot always be extended to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, as shown in the next example.
+Example 3.3.
 Let $M\coloneqq(w^{k})_{k\in\mathds{N}}\subset\mathds{R}^{2}$ with $w^{k}\coloneqq(1/2^{k},0)$, for all $k\in\mathds{N}$.
 Define the sequence $(x^{k})_{k\in\mathds{N}}\subset\mathds{R}^{2}$ by $x^{0}\coloneqq(0,2)$ and, for $\ell\in\mathds{N}$,
-$x^{2\ell+1}\coloneqq x^{2\ell}+(1/2^{\ell},0)\;,\quad x^{2\ell+2}\coloneqq\left(0,\sqrt{\|x^{2\ell+1}-(1,0)\|^{2}-1}\right).$
-If $k$ is even, $x^{k+1}$ is obtained by moving $1/2^{(k/2)}$ horizontally to the right.
-If $k$ is odd, $x^{k+1}$ is the intersection between the vertical axis and the arc centered in $w^{0}=(1,0)$ passing through $x^{k}$; see Figure 1.
-Figure 1: From Example 1: set $M$, first nine points of sequence $(x^{k})_{k\in\mathds{N}}$, $\operatorname{conv}(M)=(0,1]\times\{0\}$, $\operatorname{cl}\left({\operatorname{conv}(M)}\right)\setminus\operatorname{conv}(M)=\{(0,0)\}$ and limit point $x^{*}$ of sequence $(x^{k})_{k\in\mathds{N}}$.
-We now show that $(x^{k})_{k\in\mathds{N}}$ is Fejér* with respect to $M$.
+$x^{2\ell+1}\coloneqq x^{2\ell}+(1/2^{\ell},0)\;,\quad x^{2\ell+2}\coloneqq\left(0,\sqrt{\|x^{2\ell+1}-(1,0)\|^{2}-1}\right).$ (12)
+If $k$ is even, $x^{k+1}$ is obtained by moving $x^{k}$ horizontally to the right by an amount of $1/2^{(k/2)}$.
+If $k$ is odd, $x^{k+1}$ is the intersection between the vertical axis and the circle centered at $w^{0}=(1,0)$ passing through $x^{k}$.
+Figure˜1, illustrates the set $M$, the first nine points of sequence $(x^{k})_{k\in\mathds{N}}$, the convex hull $\operatorname{conv}(M)$, its closure $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$ and the limit point $x^{*}$ of sequence $(x^{k})_{k\in\mathds{N}}$.
+The sequence that appears in Figure˜1 as well as all the calculations and the plot itself were coded and generated using Julia language [Bezanson:2017], on a MacbookAir M4, with 24GB RAM and 10-core CPU.
+Figure 1: From Example˜3.3: set $M$, first nine points of sequence $(x^{k})_{k\in\mathds{N}}$, $\operatorname{conv}(M)=(0,1]\times\{0\}$, $\operatorname{cl}\left({\operatorname{conv}(M)}\right)\setminus\operatorname{conv}(M)=\{(0,0)\}$ and limit point $x^{*}$ of sequence $(x^{k})_{k\in\mathds{N}}$.
+We now show that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $M$.
 Denote $x^{k}=(\alpha_{k},\beta_{k})$, $k\in\mathds{N}$.
@@
 Now, consider $\lambda\in(0,1]$ and set $w\coloneqq(\lambda,0)\in\operatorname{conv}(M)$. It so happens that $\|x^{k+1}-w\|\leq\|x^{k}-w\|$, for all odd $k$. Indeed,
-$\beta_{k+1}=\sqrt{\alpha_{k}^{2}-2\alpha_{k}+\beta_{k}^{2}}\leq\sqrt{\alpha_{k}^{2}-2\alpha_{k}\lambda+\beta_{k}^{2}}\eqqcolon\gamma_{k},$
-and $\gamma_{k}$ is such that $z^{k}\coloneqq(0,\gamma_{k})$ belongs to the circumference centered in $w$ and passing through $x^{k}$.
-In other words, $z^{k}$ complies with $\|z^{k}-w\|=\|x^{k}-w\|$.
-Thus,
-$\|x^{k+1}-w\|=\sqrt{\beta_{k+1}^{2}+\lambda^{2}}\leq\sqrt{\gamma_{k}^{2}+\lambda^{2}}=\|x^{k}-w\|.$
+$\beta_{k+1}=\sqrt{\alpha_{k}^{2}-2\alpha_{k}+\beta_{k}^{2}}\leq\sqrt{\alpha_{k}^{2}-2\alpha_{k}\lambda+\beta_{k}^{2}}\eqqcolon\gamma_{k},$ (13)
+and $\gamma_{k}$ is such that $z^{k}\coloneqq(0,\gamma_{k})$ belongs to the circle centered in $w$ and passing through $x^{k}$. A simple calculation shows that $\|z^{k}-w\|=\|x^{k}-w\|$. Thus,
+$\|x^{k+1}-w\|=\sqrt{\beta_{k+1}^{2}+\lambda^{2}}\leq\sqrt{\gamma_{k}^{2}+\lambda^{2}}=\|x^{k}-w\|.$ (14)
 On the other hand, for $w^{\ell}=(1/2^{\ell},0)\in{M}$ with $\ell\geq 0$, consider $N(w^{\ell})=2\ell$.
 For all even $k\geq N(w^{\ell})$, we have $\alpha_{k}=0$, $\alpha_{k+1}\leq 1/2^{\ell}$ and $\beta_{k+1}=\beta_{k}$.
-So, $\|x^{k+1}-w^{\ell}\|\leq\|x^{k}-w^{\ell}\|$ and the sequence is Fejér* monotone w.r.t.
-$M$.
+So, $\|x^{k+1}-w^{\ell}\|\leq\|x^{k}-w^{\ell}\|$ and the sequence is Fejér* monotone with respect to $M$.
 Note that $\operatorname{conv}(M)=(0,1]\times\{0\}$ and also it holds that $\bar{w}=(0,0)\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$.
-We show now that $(x^{k})_{k\in\mathds{N}}$ is not Fejér* monotone w.r.t.
-$\{(0,0)\}$.
+We show now that $(x^{k})_{k\in\mathds{N}}$ is not Fejér* monotone with respect to $\{(0,0)\}$.
 Indeed, for $k=2\ell$, $\ell\in\mathds{N}$, we have $x^{2\ell}=(0,\beta_{2\ell})$ and $x^{{2\ell}+1}=(1/2^{\ell},\beta_{{2\ell}+1})=(1/2^{\ell},\beta_{2\ell})$.
 Thus,
-$\|x^{{2\ell}+1}-\bar{w}\|^{2}=\dfrac{1}{2^{2\ell}}+\beta_{2\ell}^{2}>\beta_{2\ell}^{2}=\|x^{2\ell}-\bar{w}\|^{2}.$
+$\|x^{{2\ell}+1}-\bar{w}\|^{2}=\dfrac{1}{2^{2\ell}}+\beta_{2\ell}^{2}>\beta_{2\ell}^{2}=\|x^{2\ell}-\bar{w}\|^{2}.$ (15)
 Finally, just for the sake of completion, we show that sequence $(x^{k})_{k\in\mathds{N}}$ converges to $x^{*}\coloneqq(0,\sqrt{4/3})$.
-In fact, for $k=2\ell$, $\ell\in\mathds{N}$, we have from Example 1 that
+In fact, for $k=2\ell$, $\ell\in\mathds{N}$, we have from Eq.˜13 that
 $\beta_{2\ell+2}^{2}$ $=\alpha_{2\ell+1}^{2}-2\alpha_{2\ell+1}+\beta_{2\ell+1}^{2}=\dfrac{1}{2^{2\ell}}-2\dfrac{1}{2^{\ell}}+\beta_{2\ell}^{2}$ (16)
 $=\displaystyle\sum\limits_{j=0}^{\ell}\dfrac{1}{2^{2j}}-2\sum\limits_{j=0}^{\ell}\dfrac{1}{2^{j}}+\beta_{0}^{2}=\displaystyle\sum_{j=0}^{\ell}\dfrac{1}{2^{2j}}\stackrel{{\scriptstyle\ell\to+\infty}}{{\to}}\dfrac{4}{3}.$ (17)
-Clearly, we have $\left\lVert x^{2\ell+1}-x^{2\ell}\right\rVert\to 0$ when $\ell\to+\infty$, so $x^{k}\to x^{*}$, as desired. ∎
-Remark 1
-One might argue that the set $M$ in Example 1 is somehow special, for $M$ has empty interior.
-However, it is possible to prove the same result for an extended example where the underlying set has nonempty interior, say, for instance, $M\coloneqq\{(w^{\prime},w^{\prime\prime})\in\mathds{R}^{2}\mid 0<w^{\prime}\leq 1,-1<w^{\prime\prime}\leq 0\}$.
-In the next proposition the closure of the convex hull of the underlying set plays a role.
-Even though a Fejér* monotone sequence w.r.t.
-$M\subset\mathcal{H}$ is not Fejér* monotone w.r.t.
-$\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, one can characterize Fejér* monotonicity in terms of the convergence of scalar sequences and inner product sequences to the closure of the convex hull.
-This result can be useful for the analysis regarding weak convergence of optimization algorithms that generate Fejér* monotone sequences.
-Proposition 2
-Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be a Fejér* monotone w.r.t. a nonempty set $M$ in $\mathcal{H}$. Then,
+Clearly, we have $\left\lVert x^{2\ell+1}-x^{2\ell}\right\rVert\to 0$ when $\ell\to+\infty$, so $x^{k}\to x^{*}$, as desired.
+Remark 3.4.
+One might argue that the set $M$ as defined in Example˜3.3 is somehow special, for $M$ has empty interior.
+However, it is possible to prove the same result for an extended example where the underlying set has nonempty interior, say, for instance, the square
+$S\coloneqq\{(w^{\prime},w^{\prime\prime})\in\mathds{R}^{2}\mid 0<w^{\prime}\leq 1,-1<w^{\prime\prime}\leq 0\},$ (18)
+instead of $M$. Note that $\operatorname{conv}(M)\subsetneq(0,1]\times(-1,0]=S$.
+We present now two counter-examples showing convex combination of Fejér* monotone sequences may fail to be Fejér* monotone, as well as the fact that composition of Fejér* monotone sequences may fail to be Fejér* monotone.
+We start with the latter.
+By composition of Fejér* monotone sequences we mean the following.
+Let $T_{1},T_{2}:\mathcal{H}\to\mathcal{H}$ be two operators that generate Fejér* sequences with respect to a target set $C\subset\mathcal{H}$, that is, sequences $(x^{k}_{i})_{k\in\mathds{N}}$ defined by $x^{k+1}_{i}=T_{i}(x^{k}_{i})$ with starting point $x^{0}_{i}\in\mathcal{H}$, for $i=1,2$, that are Fejér* monotone with respect to $C$.
+Consider the operator $T=T_{1}\circ T_{2}$ and the sequence $(x^{k})_{k\in\mathds{N}}$ defined by $x^{k+1}=T(x^{k})=T_{1}(T_{2}(x^{k}))$ with starting point $x^{0}\in\mathcal{H}$.
+The question is whether the sequence $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $C$.
+The next example shows that this is not necessarily the case.
+Example 3.5.
+Let $\mathcal{H}=\mathds{R}$ and let the target set be $C=\{0\}$. We define three auxiliary sequences for $k\in\mathds{N}$:
+$u^{k}=4^{-k},\quad v^{k}=2\cdot 4^{-k},\quad w^{k}=-4^{-(k+1)}.$ (19)
+We define the operators $T_{1},T_{2}:\mathds{R}\to\mathds{R}$ by specifying their action in the following manner:
+$T_{1}(x)$ $=\begin{cases}v^{k}&\text{if }x=u^{k},\\ 0&\text{if }x=v^{k},\\ u^{k+1}&\text{if }x=w^{k},\\ 0&\text{otherwise},\end{cases}$ $T_{2}(x)$ $=\begin{cases}u^{k}&\text{if }x=u^{k},\\ w^{k}&\text{if }x=v^{k},\\ 0&\text{if }x=w^{k},\\ 0&\text{otherwise}.\end{cases}$ (20)
+We claim now that both $T_{1}$ and $T_{2}$ generate Fejér* sequences with respect to the target set $C=\{0\}$.
+Indeed, for $T_{1}$, a trajectory starting at $u^{k}$ moves $u^{k}\to v^{k}\to 0$.
+Although $|v^{k}|>|u^{k}|$ (an expansion), the sequence hits the target in 2 steps.
+A trajectory starting at $w^{k}$ moves $w^{k}\to u^{k+1}\to v^{k+1}\to 0$.
+The remaining trajectories go directly to the target.
+In all cases, the sequence enters the target set after finitely many steps, so the tail is monotone (constant zero).
+Thus, $T_{1}$ generates Fejér* sequences.
+For $T_{2}$, a trajectory starting at $v^{k}$ moves $v^{k}\to w^{k}\to 0$.
+Since $|w^{k}|<|v^{k}|$, this is strictly contractive.
+Trajectories starting at $u^{k}$ stay fixed at $u^{k}$ (monotone).
+The remaining trajectories go directly to the target.
+Thus, $T_{2}$ also yields Fejér* sequences.
+Consider now the operator $T=T_{1}\circ T_{2}$ and the sequence $(x^{k})_{k\in\mathds{N}}$ defined by $x^{k+1}=T(x^{k})$ with starting point $x^{0}=u^{0}=1$.
+Then, the sequence $(x^{k})_{k\in\mathds{N}}$ evolves as follows:
+$u^{0}\to v^{0}\to u^{1}\to v^{1}\to u^{2}\to v^{2}\to\dots$ (21)
+Explicitly, we analyse the transitions.
+From $u^{k}$ to $v^{k}$, we have $T_{2}(u^{k})=u^{k}$, and then $T_{1}(u^{k})=v^{k}$.
+Thus, if $x^{k}=u^{k}$, then $x^{k+1}=v^{k}$.
+Here, the distance to the target $C$ changes from $|u^{k}|$ to $|v^{k}|=2|u^{k}|$, yielding $|x^{k+1}-0\ >|x^{k}-0|$, a violation of Fejér* monotonicity.
+Now, from $v^{k}$ to $u^{k+1}$ we get $T_{2}(v^{k})=w^{k}$, and then $T_{1}(w^{k})=u^{k+1}$.
+Tehrefore, if $x^{k+1}=v^{k}$, then $x^{k+2}=u^{k+1}$.
+Now, the distance changes from $|v^{k}|$ to $|u^{k+1}|=\frac{1}{8}|v^{k}|$, which corresponds to a contraction.
+This indicates that the sequence generated by $T$ oscillates indefinitely between expansion and contraction.
+Since the distance strictly increases at every step mapping $u^{k}\to v^{k}$ (which occurs infinitely often), there exists no index $K$ such that the tail of the sequence is monotone non-increasing.
+Therefore, the composition $T$ does not generate a Fejér* sequence.
+We now address the case of convex combinations of Fejér* monotone sequences.
+Example 3.6.
+Let $\mathcal{H}=\mathds{R}$ and $C=\{0\}$.
+Let $(x^{k})_{k\in\mathds{N}}$ and $(y^{k})_{k\in\mathds{N}}$ be two sequences defined by: $x^{0}=2,x^{1}=1,x^{k}=2/k$ and $y^{0}=-2,y^{1}=-1.8,y^{k}=-2+2\cdot 10^{-k}$, for $k\geq 1$.
+Both sequences are Fejér monotone with respect to $C$, since both converge to zero.
+However, consider their convex combination $w^{k}\coloneqq\frac{1}{2}x^{k}+\frac{1}{2}y^{k}$.
+We have $w^{0}=0$, which implies $\left\lvert w^{0}-0\right\rvert=0$.
+At the next step, $w^{1}=\frac{1}{2}(1-1.8)=-0.4$, yielding $\left\lvert w^{1}-0\right\rvert=0.4$.
+Note that $\left\lvert w^{1}-0\right\rvert>\left\lvert w^{0}-0\right\rvert$.
+Indeed, $w^{k}=-1+1/k+10^{-k}$, for $k\geq 1$, which is decreasing sequence converging to 1-1, yet with $\left\lvert w^{k+1}\right\rvert=\left\lvert w^{k+1}-0\right\rvert>\left\lvert w^{k}-0\right\rvert=\left\lvert w^{k}\right\rvert$.
+Thus, the sequence $(w^{k})_{k\in\mathds{N}}$ fails to be Fejér monotone with respect to $C=\{0\}$, despite being a convex combination of two Fejér sequences (converging to zero).
+Since Fejér monotone sequences are Fejér* sequences, this example also shows that convex combinations of Fejér* sequences are not necessarily Fejér* sequences.
+The next remark gives a sufficient condition for convex combinations of Fejér* monotone sequences to be Fejér* monotone.
+Remark 3.7.
+Fejér* monotonicity property is preserved under synchronized linear contraction, that is, if the sequences $(x^{k})_{k\in\mathds{N}}$ and $(y^{k})_{k\in\mathds{N}}$ approach the target $z$ along the same direction with the same step-size sequence, then any convex combination of them is also Fejér* monotone.
+Formally, let $C$ be a closed convex set and $(x^{k})_{k\in\mathds{N}}$, $(y^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be sequences.
+Suppose that there exists $z\in C$, $K_{1},K_{2}\in\mathds{N}$, and scalars $\rho_{k}\in[0,1]$ such that
+$x^{k+1}-z=\rho_{k}(x^{k}-z),\forall k\geq K_{1},\quad\text{and}\quad y^{k+1}-z=\rho_{k}(y^{k}-z),\forall k\geq K_{2}.$ (22)
+Then, any sequence $(z^{k})_{k\in\mathds{N}}$ defined as the convex combination $z^{k}=\lambda x^{k}+(1-\lambda)y^{k}$, for $\lambda\in[0,1]$, is Fejér monotone for $k\geq K\coloneqq\max(K_{1},K_{2})$, and thus constitutes a Fejér* sequence.
+In fact, for $k\geq K$, we have
+$z^{k+1}-z=\lambda\rho_{k}(x^{k}-z)+(1-\lambda)\rho_{k}(y^{k}-z)=\rho_{k}(z^{k}-z).$ (23)
+Taking norms yields $\|z^{k+1}-z\|=\rho_{k}\|z^{k}-z\|\leq\|z^{k}-z\|$, for $k\geq K$.
+This section ends with the next proposition, where the closure of the convex hull of the underlying set plays a role.
+Even though a Fejér* monotone sequence with respect to $M\subset\mathcal{H}$ is not Fejér* monotone with respect to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, one can characterize Fejér* monotonicity in terms of the convergence of scalar sequences and inner product sequences to the closure of the convex hull.
+Proposition 3.8.
+Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be Fejér* monotone with respect to a nonempty set $M$ in $\mathcal{H}$. Then,
   1. (i)
-    for every $\bar{x}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$ the scalar sequence $(\left\lVert x^{k}-\bar{x}\right\rVert)_{k\in\mathds{N}}$ converges;
+    for every $\bar{x}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$ the scalar sequence $\left(\|{x^{k}-\bar{x}}\|\right)_{k\in\mathds{N}}$ converges;
   2. (ii)
     for every $\bar{x}_{1},\bar{x}_{2}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, it holds that the sequence $\left(\left\langle\bar{x}_{1}-\bar{x}_{2},x^{k}\right\rangle\right)_{k\in\mathds{N}}$ converges.
-Proof
-Regarding item (i), recall that Proposition 1(iii) implies that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t.
-$\operatorname{conv}(M)$.
-Thus, for any $w\in\operatorname{conv}(M)$, Proposition 1(ii) yields that the scalar sequence $(\left\lVert x^{k}-w\right\rVert)_{k\in\mathds{N}}$ converges.
-Now, take $\bar{x}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, and consider a sequence $(w^{\ell})_{k\in\mathds{N}}\subset\operatorname{conv}(M)$ such that $w^{\ell}\rightarrow\bar{x}$.
+Proof.
+Regarding item (i), recall that Proposition˜3.1(iii) implies that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $\operatorname{conv}(M)$.
+Thus, for any $w\in\operatorname{conv}(M)$, Proposition˜3.1(ii) yields that the scalar sequence $(\left\lVert x^{k}-w\right\rVert)_{k\in\mathds{N}}$ converges.
+Now, take $\bar{x}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, and consider a sequence $(w^{\ell})_{\ell\in\mathds{N}}\subset\operatorname{conv}(M)$ such that $w^{\ell}\rightarrow\bar{x}$.
 Using the triangle inequality, we have, for all $\ell\in\mathds{N}$,
 $-\left\|w^{\ell}-\bar{x}\right\|\leq\left\|x^{k}-\bar{x}\right\|-\left\|x^{k}-w^{\ell}\right\|\leq\left\|w^{\ell}-\bar{x}\right\|.$ (24)
+Taking $\liminf$ and $\limsup$ with respect to $k$ in the above inequalities, we obtain
+$-\left\|w^{\ell}-\bar{x}\right\|$ $\leq\liminf_{k\to\infty}\left\|x^{k}-\bar{x}\right\|-\lim_{k\to\infty}\left\|x^{k}-w^{\ell}\right\|$ (25)
+$\leq\limsup_{k\to\infty}\left\lVert x^{k}-\bar{x}\right\rVert-\lim_{k\to\infty}\left\|x^{k}-w^{\ell}\right\|\leq\left\|w^{\ell}-\bar{x}\right\|,$ (26)
+because, as aforementioned, item Proposition˜3.1(ii) says that, for every $\ell\in\mathds{N}$, $\lim_{k\to\infty}\left\|w^{\ell}-x^{k}\right\|$ exists.
+Now, as $\ell$ goes to infinity, we have $\left\|w^{\ell}-\bar{x}\right\|\to 0$, and we conclude that
 […]
-$-\left\|w^{\ell}-\bar{x}\right\|$ $\leq\liminf_{k}\left\|x^{k}-\bar{x}\right\|-\lim_{k}\left\|x^{k}-w^{\ell}\right\|$ (7)
-$\leq\limsup_{k}\left\lVert x^{k}-\bar{x}\right\rVert-\lim_{k}\|x^{k}-w^{\ell}\|\leq\left\|w^{\ell}-\bar{x}\right\|,$ (8)
-because, as aforementioned, item Proposition 1(ii) says that, for every $\ell\in\mathds{N}$, $\lim_{k}\left\|w^{\ell}-x^{k}\right\|$ exists.
-Now, as $\ell$ goes to infinity, we have $\left\|w^{\ell}-\bar{x}\right\|\to 0$, and we conclude that
-$\liminf_{k}\left\|x^{k}-\bar{x}\right\|=\limsup_{k}\left\|x^{k}-\bar{x}\right\|=\lim_{k}\left\|x^{k}-\bar{x}\right\|,$
 so that we have the desired result.
@@
 $=\frac{1}{2}\left(\left\|x^{k}-\bar{x}_{2}\right\|^{2}-\left\|\bar{x}_{1}-\bar{x}_{2}\right\|^{2}-\left\|x^{k}-\bar{x}_{1}\right\|^{2}\right)+\left\langle\bar{x}_{1}-\bar{x}_{2},\bar{x}_{1}\right\rangle.$ (29)
-Now, item (i) implies that the sequences $(\left\|x^{k}-\bar{x}_{1}\right\|)_{k\in\mathds{N}}$ and $(\left\|x^{k}-\bar{x}_{2}\right\|)_{k\in\mathds{N}}$ converge.
+Now, item (i) implies that the sequences $\left(\left\|x^{k}-\bar{x}_{1}\right\|\right)_{k\in\mathds{N}}$ and $\left(\left\|x^{k}-\bar{x}_{2}\right\|\right)_{k\in\mathds{N}}$ converge.
 Therefore, the sequence $\left(\left\langle\bar{x}_{1}-\bar{x}_{2},x^{k}\right\rangle\right)_{k\in\mathds{N}}$ converges.
-∎
 4 Fejér* and quasi-Fejér sequences
-In this section we discuss Fejér* monotonicity relation with quasi-Fejér monotone sequences (as addressed in Combettes:2001a ).
-First, we establish the following notation: $\ell_{+}$ is the set of real sequences with nonnegative entries, and $\ell^{1}$ is the set of real sequences with finite $\ell^{1}$-norm.
-Definition 2 (quasi-Fejér monotonicity (Combettes:2001a, , Def. 1.1))
+In this section we discuss the relation between Fejér* monotonicity and quasi-Fejér monotonicity (as addressed in [Combettes:2001a]).
+We use the following notation: $\ell_{+}$ is the set of real sequences with nonnegative entries, and $\ell^{1}$ is the set of real sequences with finite $\ell_{1}$-norm.
+Definition 4.1 (quasi-Fejér monotonicity [Combettes:2001a, Def. 1.1]).
 Let $M\subset\mathcal{H}$ be a nonempty set. We say that a sequence $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$, with respect to $M$, is
@@
     quasi-Fejér of Type I if there exists a sequence $(\varepsilon_{k})_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$ such that, for all $x\in M$ and $k\in\mathds{N}$, it holds that
-    $\left\|x^{k+1}-x\right\|\leq\left\|x^{k}-x\right\|+\varepsilon_{k};$
+    $\left\|x^{k+1}-x\right\|\leq\left\|x^{k}-x\right\|+\varepsilon_{k};$ (30)
   2. (ii)
     quasi-Fejér of Type II if there exists a sequence $(\varepsilon_{k})_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$ such that, for all $x\in M$ and $k\in\mathds{N}$, it holds that
-    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}+\varepsilon_{k};$
+    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}+\varepsilon_{k};$ (31)
   3. (iii)
     quasi-Fejér of Type III if for all $x\in M$ there exists a sequence $(\varepsilon_{k})_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$ such that, for all $k\in\mathds{N}$, it holds that
-    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}+\varepsilon_{k}.$
-Sure enough, Fejér sequences are quasi-Fejér of Type I sequences.
+    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}+\varepsilon_{k}.$ (32)
+It is clear that Fejér sequences are quasi-Fejér of Type I sequences.
 Also, one can clearly see that Type II implies Type III.
-Moreover, it was established in (Combettes:2001a, , Prop.
-3.2) that if $M$ is bounded, then Type I implies Type II.
+Moreover, it was established in [Combettes:2001a, Prop.
+3.2] that if $M$ is bounded, then Type I implies Type II.
 Quasi-Fejér of Type III sequences are the most general ones.
 However, due to this generality, they are hardly used for the analysis of optimization algorithms.
-In fact, the most common type of quasi-Fejér sequences used in the literature are quasi-Fejér of Type II sequences; see, for instance, Alber:1998 ; Iusem:1994 .
-The next theorem shows that Fejér* monotonicity sequences are also quasi-Fejér of Type III.
-Theorem 4.1
-Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be a Fejér* monotone sequence w.r.t.
-a nonempty set $M$ in $\mathcal{H}$.
-Then, $(x^{k})_{k\in\mathds{N}}$ is a quasi-Fejér of Type III sequence w.r.t.
-$M$.
-Proof
-Suppose that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $M$. Then, in view of Definition 1, for any $x\in M$, there exists $N(x)$ such that, for all $k\geq N(x)$, it holds that
-$\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}.$
+In fact, the most common type of quasi-Fejér sequences used in the literature are quasi-Fejér of Type II sequences; see, for instance, [Alber:1998, Iusem:1994].
+The next theorem shows that Fejér* monotone sequences are also quasi-Fejér of Type III.
+Theorem 4.2.
+Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be a Fejér* monotone sequence with respect to a nonempty set $M$ in $\mathcal{H}$.
+Then, $(x^{k})_{k\in\mathds{N}}$ is a quasi-Fejér of Type III sequence with respect to $M$.
+Proof.
+Suppose that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $M$. Then, in view of Eq.˜2, for any $x\in M$, there exists $N(x)$ such that, for all $k\geq N(x)$, it holds that
+$\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}.$ (33)
 Let us now define the sequence $(\varepsilon_{k})_{k\in\mathds{N}}$. For $k\geq N(x)$, we set $\varepsilon_{k}\coloneqq 0$. On the other hand, for $k<N(x)$, we might have
-$\left\|x^{k+1}-x\right\|>\left\|x^{k}-x\right\|,$
+$\left\|x^{k+1}-x\right\|>\left\|x^{k}-x\right\|,$ (34)
 so we define
-$\varepsilon_{k}\coloneqq\begin{cases*}\left\|x^{k+1}-x\right\|^{2}-\left\|x^{k}-x\right\|^{2},&\text{if $\left\|x^{k+1}-x\right\|>\left\|x^{k}-x\right\|$},\\ 0,&\text{otherwise}.\end{cases*}$
-Thus, it holds true that sequence $(\varepsilon_{k})_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$ and therefore sequence $(x^{k})_{k\in\mathds{N}}$ is quasi-Fejér of Type III w.r.t. $M$. ∎
+$\varepsilon_{k}\coloneqq\begin{cases*}\left\|x^{k+1}-x\right\|^{2}-\left\|x^{k}-x\right\|^{2},&\text{if $\left\|x^{k+1}-x\right\|>\left\|x^{k}-x\right\|$},\\ 0,&\text{otherwise}.\end{cases*}$ (35)
+Thus, we conclude that $(\varepsilon_{k})_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$ and, therefore, the sequence $(x^{k})_{k\in\mathds{N}}$ is quasi-Fejér of Type III with respect to $M$. ∎
+Remark 4.3 (Fejér* versus quasi-Fejér sequences).
 We note that the concepts of Fejér* and quasi-Fejér of Type III sequences clearly are not equivalent, with the latter being more general than the former.
-Actually, if $M$ is a singleton, then any Fejér* monotone sequence is quasi-Fejér of Type II, using the same arguments used in the proof of Theorem 4.1.
+Moreover, if $M$ is a singleton, then any Fejér* monotone sequence is quasi-Fejér of Type II, using the same arguments used in the proof of Theorem˜4.2.
 Of course, in this case one can easily build quasi-Fejér sequences of Type II that are not Fejér*.
-Nevertheless, if a sequence is quasi-Fejér of Type II w.r.t.
-a general $M$, then it is Type II with respect to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$ (see (Combettes:2001a, , Prop.
-3.4)).
-Hence, in view of Example 1, we have a Fejér* sequence that it is not quasi-Fejér of Type II monotone.
-Note that we can construct a non-singleton set $M$, for instance, $M\coloneqq[-1,0]\times\{0\}\subset\mathds{R}^{2}$, such that the same sequence $(x^{k})_{k\in\mathds{N}}$ of Example 1 is quasi-Fejér of Type II, but not Fejér* monotone, w.r.t.
-$M$.
-More than not being equivalent, we have just seen that the concepts of Fejér* and quasi-Fejér of Type II are not contained in each other, that is, those ideas are independent.
-The previous theorem reveals a particular quasi-Fejér of type III sequence that is related to optimization algorithms, through Fejér* monotonicity Behling:2024c .
+Nevertheless, if a sequence is quasi-Fejér of Type II with respect to a general nonemtpy $M$, then it is Type II with respect to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$ (see [Combettes:2001a, Prop.
+3.4]).
+Hence, in view of Example˜3.3, we have a Fejér* sequence that is not quasi-Fejér of Type II.
+Note that we can construct a non-singleton set $M$, for instance, $M\coloneqq[-1,0]\times\{0\}\subset\mathds{R}^{2}$, such that the same sequence $(x^{k})_{k\in\mathds{N}}$ of Example˜3.3 is quasi-Fejér of Type II, but not Fejér* monotone with respect to $M$.
+Therefore, more than not being equivalent, we have just seen that the concepts of Fejér* and quasi-Fejér of Type II do not imply each other, that is to say, these concepts are independent.
+The previous theorem reveals a particular quasi-Fejér of Type III sequence that is related to optimization algorithms (as seen in Section˜2), through Fejér* monotonicity.
 Thus, we end our discussion by presenting some characterizations of Fejér* monotonicity in terms of the weak convergence of the underlying sequence.
-These statements are inspired by Combettes:2001a and could be derived from the results therein, since Fejér* monotonicity implies quasi-Fejér of type III monotonicity.
-However, we present them here, for the sake of completeness.
-Moreover, the provided proofs are simpler and more clear for the reader, in the context of Fejér* monotonicity.
+These statements are inspired by [Combettes:2001a] and could be derived from the results therein, since Fejér* monotonicity implies quasi-Fejér of Type III monotonicity.
 From now on, for a sequence $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$, we denote by $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$ the set of its weak cluster points and by $\mathfrak{S}\left(x^{k}\right)_{k\in\mathds{N}}$ the set of its strong cluster points.
+We remind that for a nonempty set $M$, the affine hull of $M$ is $\operatorname{aff}(M)\coloneqq\left\{\sum_{i=1}^{p}\lambda_{i}x_{i}\mid x_{i}\in M,\,\lambda_{i}\in\mathds{R},\,\sum_{i=1}^{p}\lambda_{i}=1,p\in\mathds{N}\right\},$ that is, the smallest affine subspace containing $M$.
 First, we give portrayals of Fejér* monotonicity in terms of the weak convergence of the sequence.
-Proposition 3
-Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to $M\subset\mathcal{H}$. Then,
+Proposition 4.4.
+Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be Fejér* sequence with respect to nonempty $M\subset\mathcal{H}$. Then,
   1. (i)
@@
     for any $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$, there exists an $\alpha\in\mathds{R}$ such that
-    $M\subset\left\{y\in\mathcal{H}\mid\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha\right\};$
+    $M\subset\left\{y\in\mathcal{H}\mid\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha\right\};$ (36)
   3. (iii)
@@
     if $x^{k}\rightharpoonup\bar{x}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, then the scalar sequence $\left(\left\|x^{k}-y\right\|\right)_{k\in\mathds{N}}$ converges for every $y\in\mathcal{H}$.
-Proof
-Item (i): the statement is valid since, due to Proposition 1(i), sequence $\left(x^{k}\right)_{k\in\mathds{N}}$ is bounded.
+Proof.
+Item (i): the statement is valid since, due to Proposition˜3.1(i), the sequence $\left(x^{k}\right)_{k\in\mathds{N}}$ is bounded.
 Item (ii): Take any $y\in M$. For all $k\in\mathds{N}$, we get
-$\left\|x^{k}-y\right\|^{2}-\|y\|^{2}=\left\|x^{k}\right\|^{2}-2\left\langle y,x^{k}\right\rangle.$
-Moreover, it follows from Proposition 1(ii) that $\lim\left(\left\|x^{k}-y\right\|^{2}-\|y\|^{2}\right)$ is well-defined.
-Now, given two points $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$, such that $x^{j_{k}}\rightharpoonup w^{\prime}$ and $x^{\ell_{k}}\rightharpoonup w^{\prime\prime}$, and considering Section 4 we have
-$\lim\left\|x^{j_{k}}\right\|^{2}-2\left\langle y,w^{\prime}\right\rangle=\lim\left\|x^{\ell_{k}}\right\|^{2}-2\left\langle y,w^{\prime\prime}\right\rangle,$
+$\left\|x^{k}-y\right\|^{2}-\|y\|^{2}=\left\|x^{k}\right\|^{2}-2\left\langle y,x^{k}\right\rangle.$ (37)
+Moreover, it follows from Proposition˜3.1(ii) that $\lim_{k\to\infty}\left(\left\|x^{k}-y\right\|^{2}-\|y\|^{2}\right)$ exists.
+Now, given two points $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$, such that $x^{j_{k}}\rightharpoonup w^{\prime}$ and $x^{\ell_{k}}\rightharpoonup w^{\prime\prime}$, and considering Eq.˜37, we have
+$\lim_{k\to\infty}\left\|x^{j_{k}}\right\|^{2}-2\left\langle y,w^{\prime}\right\rangle=\lim_{k\to\infty}\left\|x^{\ell_{k}}\right\|^{2}-2\left\langle y,w^{\prime\prime}\right\rangle,$ (38)
 and therefore
-$\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\frac{1}{2}\left(\lim\left\|x^{j_{k}}\right\|^{2}-\lim\left\|x^{\ell_{k}}\right\|^{2}\right).$
+$\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\frac{1}{2}\left(\lim_{k\to\infty}\left\|x^{j_{k}}\right\|^{2}-\lim_{k\to\infty}\left\|x^{\ell_{k}}\right\|^{2}\right).$ (39)
 Setting $\alpha$ as the right-hand side of the above equation yields that
-$M\subset\left\{y\in\mathcal{H}\mid\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha\right\},$
+$M\subset\left\{y\in\mathcal{H}\mid\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha\right\},$ (40)
 as required.
@@
 Consequently, $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$ reduces to a singleton.
-Due to Proposition 2(i), $\left(x^{k}\right)_{k\in\mathds{N}}$ lies in a weakly compact set and therefore converges weakly.
+Due to Proposition˜3.8(i), $\left(x^{k}\right)_{k\in\mathds{N}}$ lies in a weakly compact set and therefore converges weakly.
 Item (iv): Consider any $y\in\mathcal{H}$. Then, for all $k\in\mathds{N}$, we have
-$\left\|x^{k}-y\right\|^{2}=\left\|x^{k}-\bar{x}\right\|^{2}+2\left\langle x^{k}-\bar{x},\bar{x}-y\right\rangle+\|\bar{x}-y\|^{2}.$
-Using the hypothesis and Proposition 2 on the right-hand side of the above equation, we conclude the convergence of $\left(\left\|x^{k}-y\right\|\right)_{k\in\mathds{N}}$.
-Theorem 4.2
-Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to a nonempty set $M\subset\mathcal{H}$.
+$\left\|x^{k}-y\right\|^{2}=\left\|x^{k}-\bar{x}\right\|^{2}+2\left\langle x^{k}-\bar{x},\bar{x}-y\right\rangle+\|\bar{x}-y\|^{2}.$ (41)
+Using the hypothesis and Proposition˜3.8 on the right-hand side of the above equation, we conclude the convergence of $\left(\left\|x^{k}-y\right\|\right)_{k\in\mathds{N}}$.
+Theorem 4.5.
+Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a sequence in $\mathcal{H}$ and let $M$ be a nonempty subset of $\mathcal{H}$.
+Suppose that $\left(x^{k}\right)_{k\in\mathds{N}}$ is Fejer* monotone with respect to $M$.
 Then, $\left(x^{k}\right)_{k\in\mathds{N}}$ converges weakly to a point in $M$ if, and only if, $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\subset M$.
-Proof
-It holds trivially that Fejér* monotonicity together with weakly convergence implies that $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\subset M$.
-On the other hand, suppose that $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\subset M$ and take $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$.
-Thus, $w^{\prime},w^{\prime\prime}\in M$.
-Proposition 3(ii) affirms that $\left\langle w^{\prime},w^{\prime}-w^{\prime\prime}\right\rangle=\left\langle w^{\prime\prime},w^{\prime}-w^{\prime\prime}\right\rangle$ and hence, $w^{\prime}=w^{\prime\prime}$.
-The boundedness of $\left(x^{k}\right)_{k\in\mathds{N}}$, guaranteed by Proposition 1, completes the proof.
-∎
+Proof.
+The result follows directly from Theorem˜4.2 and [Combettes:2001a, Thm. 3.8]. ∎
 We conclude this section by presenting a pair of results that characterizes the strong convergence of Fejér* monotone sequences.
-Proposition 4
-Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to a nonempty set $M\subset\mathcal{H}$. Then,
+Proposition 4.6.
+Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a sequence in $\mathcal{H}$ and let $M$ be a nonempty subset of $\mathcal{H}$.
+Suppose that $\left(x^{k}\right)_{k\in\mathds{N}}$ is Fejer* monotone with respect to $M$.
+Then,
   1. (i)
     for all $w^{\prime},w^{\prime\prime}\in\mathfrak{S}\left(x^{k}\right)_{k\in\mathds{N}}$, it holds that
-    $M\subset\left\{y\in\mathcal{H}\mid\left\langle y-\frac{w^{\prime}+w^{\prime\prime}}{2},w^{\prime}-w^{\prime\prime}\right\rangle=0\right\};$
+    $M\subset\left\{y\in\mathcal{H}\mid\left\langle y-\frac{w^{\prime}+w^{\prime\prime}}{2},w^{\prime}-w^{\prime\prime}\right\rangle=0\right\};$ (42)
   2. (ii)
@@
     the sequence converges strongly if there exist $x\in M,\left(\varepsilon_{k}\right)_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$, and $\rho\in(0,+\infty)$ such that for all $k\in\mathds{N}$ it holds that
-    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}-\rho\left\|x^{k+1}-x^{k}\right\|+\varepsilon_{k}.$
-Proof
-It follows from Theorem 4.1 together with Proposition 3.7 in Combettes:2001a . ∎
-Theorem 4.3
-Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to a nonempty set $M\subset\mathcal{H}$. Then, the following statements are equivalent:
+    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}-\rho\left\|x^{k+1}-x^{k}\right\|+\varepsilon_{k}.$ (43)
+Proof.
+It follows from Theorem˜4.2 together with Proposition 3.9 in [Combettes:2001a]. ∎
+Theorem 4.7.
+Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a sequence in $\mathcal{H}$ and let with $M$ be a nonempty subset of $\mathcal{H}$.
+Suppose that $\left(x^{k}\right)_{k\in\mathds{N}}$ is Fejer* monotone with respect to $M$.
+Then, the following statements are equivalent:
   1. (i)
@@
     $\mathfrak{S}\left(x^{k}\right)_{k\in\mathds{N}}\cap M\neq\varnothing$.
-Proof
-This is a direct result of combining Theorem 4.1 and Theorem 3.11 in Combettes:2001a . ∎
-We emphasize that new characterizations of Fejér* monotonicity could be revealed with further analysis, suggesting a direction for future research.
+Proof.
+This is a direct result of combining Theorem˜4.2 and Theorem 3.11 in [Combettes:2001a].
 5 Concluding remarks
-In this work we further expand the understanding of Fejér* monotonicity.
-We provide a characterization of Fejér* monotonicity in terms of the convergence of scalar sequences and inner product sequences to the closure of the convex hull of the underlying set.
-We provide an insightful example that shows that Fejér* monotonicity not always can be extended to the closure of the convex hull of the underlying set.
-It was also established connections between Fejér* monotonicity and quasi-Fejér monotonicity.
-The results presented in this paper could shed a light on the analysis of the (weak) convergence of algorithms that comply with Fejér* monotonicity.
-Acknowledgements.
-RB was partially supported by Brazilian agency Fundação de Amparo à Pesquisa do Estado do Rio de Janeiro (Grant E-26/201.345/2021); YBC was partially supported by the USA agency National Science Foundation (Grant DMS-2307328); AAR was partially supported by Brazilian agency Conselho Nacional de Desenvolvimento Científico e Tecnológico – CNPq (Grant 307987/2023-0); LRS was also partially funded by CNPq (Grant 310571/2023-5).
+In this work, we further expanded the understanding of Fejér* monotonicity.
+We provided a characterization of Fejér* monotonicity in terms of the convergence of scalar sequences and inner product sequences to the closure of the convex hull of the underlying set.
+The insightful Example˜3.3 illustrates that Fejér* monotonicity cannot, in general, be extended to the closure of the convex hull of the underlying set.
+We also established connections between Fejér* monotonicity and quasi-Fejér monotonicity, highlighting that Fejér* monotonicity lies outside the classical Fejér and quasi-Fejér of Type II frameworks.
+Fejér* monotonicity notion has already been used to establish convergence (usually finite) of algorithms in [Iusem:1986a, fukushimaFinitelyConvergentAlgorithm1982, iusemFinitelyConvergentIterative1987, DePierro:1988a, Kolobov:2022] and, more recently, was formally defined in [Behling:2024c].
+The results presented in this paper could be used to shed light on the analysis of the (weak) convergence of algorithms that comply with Fejér* monotonicity, mainly to aid to prove finite convergence results.
+[…]
+Acknowledgements The authors would like to thank the anonymous reviewers for their careful reading of our manuscript and their insightful comments and suggestions that helped us to improve the quality of this paper.
+In particular, the discussions presentend in Examples˜3.5, 3.6, and 3.7 were raised by one of the reviewers.
+RB was partially supported by Brazilian agency Fundação de Amparo à Pesquisa do Estado do Rio de Janeiro (Grant E-26/201.345/2021); YBC was partially supported by the USA agency National Science Foundation (Grant DMS-2307328); AAR was partially supported by Brazilian agency Conselho Nacional de Desenvolvimento Científico e Tecnológico – CNPq (Grant 307987/2023-0); LRS was also partially funded by Brazilian agencies CNPq (Grants 310571/2023-5 and 407147/2023-3) and Fundação de Amparo ao Estado de Santa Catatrina – FAPESC (Edital 21/2024, Grant 2024TR002238).
+[…]
+AI use declaration During the preparation of this work, the authors used Google Gemini (Large Language Model) for language editing, generating code, and refining the clarity of the text.
+The authors reviewed and verified all AI-generated content and take full responsibility for the accuracy and integrity of the publication.